### v2

# Geometric structure of shallow neural networks and constructive ${\mathcal{L}}^{2}$ cost minimization

Abstract.
In this paper, we approach the problem of cost (loss) minimization in underparametrized shallow neural networks through the explicit construction of upper bounds, without any use of gradient descent. A key focus is on elucidating the geometric structure of approximate and precise minimizers. We consider shallow neural networks with one hidden layer, a ReLU activation function, an ${\mathcal{L}}^{2}$ Schatten class (or Hilbert-Schmidt) cost function, input space ${\mathbb{R}}^{M}$, output space ${\mathbb{R}}^{Q}$ with $Q\leq M$, and training input sample size $N>QM$ that can be arbitrarily large. We prove an upper bound on the minimum of the cost function of order $O(\delta_{P})$ where $\delta_{P}$ measures the signal to noise ratio of training inputs. In the special case $M=Q$, we explicitly determine an exact degenerate local minimum of the cost function, and show that the sharp value differs from the upper bound obtained for $Q\leq M$ by a relative error $O(\delta_{P}^{2})$. The proof of the upper bound yields a constructively trained network; we show that it metrizes a particular $Q$-dimensional subspace in the input space ${\mathbb{R}}^{M}$. We comment on the characterization of the global minimum of the cost function in the given context.
1. Introduction
Applications of shallow and deep neural networks have rapidly become indispensable in a vast range of research disciplines, and their technological impact in recent years has been enormous. However, despite of those successes, the fundamental conceptual reasons underlying their functioning are currently insufficiently understood, and remain the subject of intense investigation, [2, 6]. The common approach to the minimization of the cost (loss) function is based on the gradient descent flow generated by it, and a variety of powerful algorithms have been developed for this purpose that can be highly successful in suitable application scenarios; but nevertheless, even if a global or sufficiently good local minimum is obtained, the properties of the minimizing parameters (weights, biases) remain quite mysterious.
In this paper, we address the cost minimization problem via explicit construction of upper bounds on the global minimum of the cost function, without invoking gradient descent at all. The inspiration for our approach is drawn from methods in mathematical physics aimed at determining the ground state energy for complex manybody quantum systems (such as large molecules), see for instance [9]. Our main goal here is to obtain a rigorous mathematical understanding of the geometric structure of (approximate) cost minimizers; we do not focus on the development of efficient algorithms, or exploration of application cases in the present work. To make our exposition broadly accessible, we will mostly employ standard mathematical terminology, while minimally referring to specialized nomenclatures specific to the neural networks literature.
The current paper is the first in a series of works in which we investigate the geometric structure of cost minimizers in neural networks (that is, weights and biases that minimize the cost). Here, we analyze underparametrized shallow neural networks in the context of supervised learning, with one hidden layer, a ReLU activation function, an ${\mathcal{L}}^{2}$ Schatten class (or Hilbert-Schmidt) cost function, input space ${\mathbb{R}}^{M}$, output space ${\mathbb{R}}^{Q}$ with $Q\leq M$, and training input sample size $N>QM$ that can be arbitrarily large - therefore, we are considering the underparametrized regime. We prove an upper bound on the minimum of the cost function of order $O(\delta_{P})$ where $\delta_{P}$ measures the signal to noise ratio of training inputs. First, we derive an upper bound on the minimum of the cost function with an explicit construction of suitable weights and biases. As a by-product of our proof, we obtain a constructively trained shallow network with explicit parameters. In the special case $M=Q$, we explicitly determine an exact degenerate local minimum of the cost function, and show that the sharp value differs from the upper bound obtained for $Q\leq M$ by a relative error $O(\delta_{P}^{2})$.
The above assumptions allow for a succinct exposition of core geometric aspects of the problem; an analysis of more general situations is left for future work. For some thematically related background, see for instance [1, 7, 6, 8, 10, 11] and the references therein. We present the explicit construction of local and global ${\mathcal{L}}^{2}$ cost minimizers in underparametrized Deep Learning (that is, multilayer) networks in a separate work, [4], which will invoke results developed in the paper at hand, combined with inspiration drawn from rigorous renormalization group analysis in quantum field theory, as for instance in [3].
In this introductory section, we summarize the main results of this paper. We consider a shallow network for which $y_{j}\in{\mathbb{R}}^{Q}$ denotes the $j$-th output vector, and define the output matrix
$Y:=[y_{1},\dots,y_{Q}]\in{\mathbb{R}}^{Q\times Q}\,,$ (1.1)
which we assume to be invertible. We assume the input space to be given by ${\mathbb{R}}^{M}$ with $M\geq Q$, and for $j=1,\dots,Q$, we let
$X_{0,j}:=[x_{0,j,1}\cdots x_{0,j,i}\cdots x_{0,j,N_{j}}]\in{\mathbb{R}}^{M\times N_{j}}\,.$ (1.2)
denote the matrix of training input vectors $x_{0,j,1}$, $i=1,\dots,N_{j}$, which belong to the output $y_{j}$. Letting $N:=\sum_{j=1}^{Q}N_{j}$, the full matrix of training inputs is given by
$X_{0}:=[X_{0,1}\cdots X_{0,j}\cdots X_{0,Q}]\in{\mathbb{R}}^{M\times N}\,.$ (1.3)
We assume that $Q\leq M\leq MQ<N$ where $N$ can be arbitrarily large. We define the average of all training input vectors belonging to the output $y_{j}$,
$\overline{x_{0,j}}:=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}x_{0,j,i}\in{\mathbb{R}}^{M}$ (1.4)
for $j=1,\dots,Q$, and
$\Delta x_{0,j,i}:=x_{0,j,i}-\overline{x_{0,j}}\,.$ (1.5)
Moreover, we let
$\delta:=\sup_{i,j}|\Delta x_{0,j,i}|\,.$ (1.6)
We define
$\Delta X_{0,j}:=[\Delta x_{0,j,1}\cdots\Delta x_{0,j,i}\cdots\Delta x_{0,j,N_{j}}]\in{\mathbb{R}}^{M\times N_{j}}\,,$ (1.7)
and
$\Delta X_{0}:=[\Delta X_{0,1}\cdots\Delta X_{0,j}\cdots\Delta X_{0,Q}]\in{\mathbb{R}}^{M\times N}\,.$ (1.8)
Moreover, we define
$\overline{X_{0}^{red}}:=[\overline{x_{0,1}}\cdots\overline{x_{0,Q}}]\in{\mathbb{R}}^{M\times Q}$ (1.9)
where the superscript indicates that this is the column-wise reduction of the matrix $\overline{X_{0}}$ in (3.4), below.
We may then define the orthoprojector $P=P^{2}=P^{T}$ onto the span of the family $\{\overline{x_{0,j}}\}_{j=1}^{Q}$ which we assume to be linearly independent. We let $P^{\perp}:={\bf 1}-P$.
We consider a shallow network with hidden layer
$x\mapsto\sigma(W_{1}x+b_{1})\,,$ (1.10)
determined by a weight matrix $W_{1}\in{\mathbb{R}}^{M\times M}$, bias vector $b_{1}\in{\mathbb{R}}^{M}$, and ReLU activation function $\sigma$, which acts component-wise as a ramp function $(a)_{+}:=\max\{0,a\}$. We note that the property of $\sigma$ acting component-wise singles out the given coordinate system in ${\mathbb{R}}^{M}$ as distinct.
We define the output layer with the map ${\mathbb{R}}^{M}\rightarrow{\mathbb{R}}^{Q}$ given by
$x^{\prime}\mapsto W_{2}x^{\prime}+b_{2}$ (1.11)
where $W_{2}\in{\mathbb{R}}^{Q\times M}$ and $b_{2}\in{\mathbb{R}}^{Q}$. Here, we do not include an activation function.
Then, the ${\mathcal{L}}^{2}$ cost function ${\mathcal{C}}[W_{j},b_{j}]$ is defined as
${\mathcal{C}}[W_{j},b_{j}]:=\sqrt{\frac{1}{N}\sum_{j=1}^{Q}\sum_{i=1}^{N_{j}}|W_{2}\sigma(W_{1}x_{0,j,i}+b_{1})+b_{2}-y_{j}|^{2}_{{\mathbb{R}}^{Q}}}\,$ (1.12)
where $|\cdot|_{{\mathbb{R}}^{Q}}$ denotes the Euclidean norm on ${\mathbb{R}}^{Q}$.
In Theorem 3.1, we prove an upper bound on the minimum of the cost function with the following construction.
Let $R\in O(M)$ denote an orthogonal matrix that diagonalizes $P$. That is, the orthoprojectors
$P_{R}:=RPR^{T}\;\;\;{\rm and}\;P^{\perp}_{R}:=RP^{\perp}R^{T}$ (1.13)
are diagonal in the given coordinate system; this is important for compatibility with the fact that $\sigma$ acts component-wise. Namely, $R$ rotates the input data in such a way that ${\rm range}(P)$ is made to align with the coordinate axes; this decouples the action of $\sigma$ on the rotated ${\rm range}(P)$ from its action on the rotated ${\rm range}(P^{\perp})$. That is, $\sigma(v)=\sigma(P_{R}v)+\sigma(P^{\perp}_{R}v)$ for all $v\in{\mathbb{R}}^{M}$.
Letting
$u_{M}:=(1,1,\cdots,1)^{T}\in{\mathbb{R}}^{M}$ (1.14)
we may choose $\beta_{1}\geq 0$ sufficiently large ($\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|$ is sufficient), so that the projected, translated, and rotated training input vectors $RPx_{0,j,i}+\beta_{1}u_{M}$ are component-wise non-negative for all $i=1,\dots,N_{j}$, and $j=1,\dots,Q$.
Then, we construct an upper bound on the cost function by use of the following weights and biases.
We choose
$W_{1}^{*}=R$ (1.15)
and $b_{1}^{*}=P_{R}b_{1}^{*}+P_{R}^{\perp}b_{1}^{*}$ where
$P_{R}b_{1}^{*}=\beta_{1}P_{R}u_{Q}\;\;\;,\;\;\;P^{\perp}_{R}b_{1}^{*}=-\delta P^{\perp}_{R}u_{M}\,,$ (1.16)
with $\beta_{1}\geq 0$ large enough to ensure component-wise non-negativity of $RPX_{0}+P_{R}B_{1}^{*}$, where $B_{1}^{*}=b_{1}^{*}u_{N}^{T}$, so that
$\sigma(RPX_{0}+P_{R}B_{1}^{*})=RPX_{0}+P_{R}B_{1}^{*}\,.$ (1.17)
On the other hand, the $M-Q$ rows of $RP^{\perp}X_{0}+P_{R}^{\perp}B_{1}^{*}$ are eliminated by way of
$\sigma(RP^{\perp}X_{0}+P_{R}^{\perp}B_{1}^{*})=0\,.$ (1.18)
We thus obtain a reduction of the dimension of the input space from $M$ to $Q$.
Passing to the output layer, we require $W_{2}^{*}$ to solve
$W_{2}^{*}R\overline{X_{0}^{red}}=Y\,,$ (1.19)
which yields
$W_{2}^{*}=Y\;{\rm Pen}[\overline{X_{0}^{red}}]PR^{T}\,.$ (1.20)
Here,
${\rm Pen}[\overline{X_{0}^{red}}]:=((\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}})^{-1}(\overline{X_{0}^{red}})^{T}$ (1.21)
denotes the Penrose inverse of $\overline{X_{0}^{red}}$, which satisfies
${\rm Pen}[\overline{X_{0}^{red}}]\overline{X_{0}^{red}}={\bf 1}_{Q\times Q}\;\;\;,\;\;\overline{X_{0}^{red}}{\rm Pen}[\overline{X_{0}^{red}}]=P\,.$ (1.22)
Finally, we find that
$b_{2}^{*}=-W_{2}^{*}P_{R}b_{1}^{*}\,,$ (1.23)
that is, it reverts the translation by $P_{R}b_{1}^{*}$ in the previous layer.
A geometrically intuitive way of thinking about this construction is that $W_{1}^{*}=R$ orients the training input data with respect to the given coordinate system, in order to align it with the component-wise action of $\sigma$. This allows for a maximal rank reduction via $\sigma$, whereby the maximal possible amount of insignificant information is eliminated: $P_{R}b_{1}^{*}$ pulls the significant information (in the range of $P_{R}$) out of the kernel of $\sigma$, while $P_{R}^{\perp}b_{1}^{*}$ pushes the insignificant information (in the range of $P_{R}^{\perp}$) into the kernel of $\sigma$ whereby the latter is eliminated. Subsequently, $b_{2}^{*}$ places the significant information back into its original position, and $W_{2}^{*}$ matches it to the output matrix $Y$ in the sense of least squares, with respect to the ${\mathcal{L}}^{2}$-norm.
We refer to the shallow network defined with these specific weights and biases $W_{i}^{*},b_{i}^{*}$, $i=1,2$, as the constructively trained network. The parameter
$\delta_{P}:=\sup_{j,i}\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,j,i}\right|\,,$ (1.24)
measures the relative size between $\Delta X_{0}$ and $\overline{x_{0,j}}$, $j=1,\dots,Q$; in particular, $\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,j,i}\right|$ scales like $\frac{|\Delta x|}{|x|}$, and is scaling invariant under $X_{0}\rightarrow\lambda X_{0}$. We then obtain
$\min_{W_{j},b_{j}}{\mathcal{C}}[W_{j},b_{j}]$$\leq$ ${\mathcal{C}}[W_{i}^{*},b_{i}^{*}]$ (1.25)
$\leq$ $\frac{1}{\sqrt{N}}\|Y{\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}\|_{{\mathcal{L}}^{2}}$
$\leq$ $\|Y\|_{op}\;\delta_{P}\;\,.$
In order to match an arbitrary test input $x\in{\mathbb{R}}^{M}$ with one of the output vectors $y_{j}$, $j\in\{1,\dots,Q\}$, we let for $x\in{\mathbb{R}}^{M}$,
${\mathcal{C}}_{j}[x]:=|W_{2}^{*}\sigma(W_{1}^{*}x+b_{1}^{*})+b_{2}^{*}-y_{j}|$ (1.26)
with the weights and biases of the constructively trained network. Then,
$j^{*}={\rm argmin}_{j}{\mathcal{C}}_{j}[x]$ (1.27)
implies that $x$ matches the $j^{*}$-th output $y_{j^{*}}$.
We define the metric
$d_{\widetilde{W}_{2}}(x,y):=|\widetilde{W}_{2}P(x-y)|_{{\mathbb{R}}^{Q}}\;\;\;{\rm for}\;x,y\in{\rm range}(P)\subset{\mathbb{R}}^{M}$ (1.28)
where $\widetilde{W}_{2}:=W_{2}^{*}R$, on the $Q$-dimensional linear subspace ${\rm range}(P)\subset{\mathbb{R}}^{M}$. In Theorem 3.3, we prove that
${\mathcal{C}}_{j}[x]=d_{\widetilde{W}_{2}}(Px,x_{0,j})\,.$ (1.29)
Therefore, matching an input $x\in{\mathbb{R}}^{M}$ with an output $y_{j^{*}}$ via the constructively trained network is equivalent to solving the metric minimization problem
$j^{*}={\rm argmin}_{j\in\{1,\dots,Q\}}(d_{\widetilde{W}_{2}}(Px,\overline{x_{0,j}}))$ (1.30)
on the range of $P$.
In other words, $x\in{\mathbb{R}}^{M}$ is matched with the output $y_{j^{*}}\in{\mathbb{R}}^{Q}$ by determining that $Px$ is closest to $\overline{x_{0,j^{*}}}$ among all $\{\overline{x_{0,j}}\in{\rm range}(P)\subset{\mathbb{R}}^{M}\;|\;j=1,\dots,Q\}$ in the $d_{\widetilde{W}_{2}}$ metric.
In Theorem 3.2, we focus on the special case $M=Q$, and present the explicit construction of a degenerate local minimum of the cost function (more precisely, a weighted version of the cost function), and obtain an improvement on the upper bound (1.25) by a factor $1-C_{0}\delta_{P}^{2}$ for some constant $C_{0}\geq 0$. Here, degeneracy means that for all weights and biases $W_{1},b_{1}$ satisfying (1.17), the corresponding minimum of the cost function attains the same value. In particular, this implies that in this range of weights and biases, (1.25) differs from the sharp value with a relative error of order $O(\delta_{P}^{2})$, when $Q=M$.
In Theorem 3.5, we prove a result for $Q=M$ closely related to Theorem 3.2, but for the case where (1.17) is not satisfied.
We expect our results for the case $Q=M$ to be extensible to $Q<M$, but leave a detailed analysis for future work.
This paper is organized as follows. In Section 2, we give a detailed introduction of the mathematical model describings the shallow network. In Section 3, we present the main results, namely Theorem 3.1, which we prove in Section 4, Theorem 3.3, which we prove in Section 7, and Theorem 3.5. In Section 6, we provide a detailed description of the constructively trained network.
2. Definition of the mathematical model
In this section, we give a detailed introduction of the mathematical model in discussion, consisting of a shallow network with input layer, one hidden layer, the output layer, and an ${\mathcal{L}}^{2}$ cost function. Letting $Q\in{\mathbb{N}}$ denote the number of distinct output values of the shallow network, we define the output matrix
$Y:=[y_{1},\dots,y_{Q}]\in{\mathbb{R}}^{Q\times Q}$ (2.1)
where $y_{j}\in{\mathbb{R}}^{Q}$ is the $j$-th output vector. We assume that the family $\{y_{j}\}$ is linearly independent, so that $Y$ is invertible.
We let ${\mathbb{R}}^{M}$ denote the input space with coordinate system defined by the orthonormal basis vectors $\{e_{\ell}=(0,\dots,0,1,0,\dots,0)^{T}\in{\mathbb{R}}^{M}\,|\,\ell=1,\dots,M\}$. For $j\in\{1,\dots,Q\}$, and $N_{j}\in{\mathbb{N}}$,
$x_{0,j,i}\in{\mathbb{R}}^{M}\;\;,\;\;i\in\{1,\dots,N_{j}\}$ (2.2)
denotes the $i$-th training input vector corresponding to the $j$-th output vector $y_{j}$. We define the matrix of all training inputs belonging to $y_{j}$ by
$X_{0,j}:=[x_{0,j,1}\cdots x_{0,j,i}\cdots x_{0,j,N_{j}}]\,.$ (2.3)
Correspondence of a training input vector to the same output $y_{j}$ defines an equivalence relation, that is, for each $j\in\{1,\dots,Q\}$ we have $x_{0,j,i}\sim x_{0,j,i^{\prime}}$ for any $i,i^{\prime}\in\{1,\dots,N_{j}\}$. Accordingly, $X_{0,j}$ labels the equivalence class of all inputs belonging to the $j$-th output $y_{j}$.
Then, we define the matrix of training inputs
$X_{0}:=[X_{0,1}\cdots X_{0,j}\cdots X_{0,Q}]\in{\mathbb{R}}^{M\times N}$ (2.4)
where $N:=\sum_{j=1}^{Q}N_{j}$.
We introduce weight matrices
$W_{1}\in{\mathbb{R}}^{M\times M}\;\;,\;\;W_{2}\in{\mathbb{R}}^{Q\times M}$ (2.5)
and bias matrices
$B_{1}$ ::=:= $b_{1}u_{N}^{T}=[b_{1}\cdots b_{1}]\in{\mathbb{R}}^{M\times N}$
$B_{2}$ ::=:= $b_{2}u_{N}^{T}=[b_{2}\cdots b_{2}]\in{\mathbb{R}}^{Q\times N}$ (2.6)
where $b_{1}\in{\mathbb{R}}^{M}$ and $b_{2}\in{\mathbb{R}}^{Q}$ are column vectors, and
$u_{N}=(1,1,\dots,1)^{T}\;\;\in{\mathbb{R}}^{N}\,.$ (2.7)
Moreover, we choose to consider a ReLU activation function
$\sigma:{\mathbb{R}}^{M\times N}$$\rightarrow$ ${\mathbb{R}}_{+}^{M\times N}$
$A=[a_{ij}]$ $\mapsto$ $[(a_{ij})_{+}]$ (2.8)
where
$(a)_{+}:=\max\{0,a\}$ (2.9)
is the standard ramp function.
We define
$X^{(1)}:=\sigma(W_{1}X_{0}+B_{1})\in{\mathbb{R}}_{+}^{M\times N}$ (2.10)
for the hidden layer, and
$X^{(2)}:=W_{2}X^{(1)}+B_{2}\in{\mathbb{R}}^{Q\times N}\,$ (2.11)
for the terminal layer.
We also define the extension of the matrix of outputs, $Y^{ext}$, by
$Y^{ext}:=[Y_{1}\cdots Y_{Q}]\in{\mathbb{R}}^{Q\times N}$ (2.12)
where
$Y_{j}:=[y_{j}\cdots y_{j}]\in{\mathbb{R}}^{Q\times N_{j}}$ (2.13)
with $N_{j}$ copies of the same output column vector $y_{j}$. Clearly, $Y^{ext}$ has full rank $Q$.
Then, we consider the ${\mathcal{L}}^{2}$ Schatten class (or Hilbert-Schmidt) cost function
${\mathcal{C}}[W_{j},b_{j}]:=\frac{1}{\sqrt{N}}\|X^{(2)}-Y^{ext}\|_{{\mathcal{L}}^{2}}\,$ (2.14)
with
$\|A\|_{{\mathcal{L}}^{2}}\equiv\sqrt{{\rm Tr}(AA^{T})}\,,$ (2.15)
where $A^{T}$ is the transpose of the matrix $A$. ${\mathcal{C}}[W_{j},b_{j}]$ is a function of the weights and biases, while the training inputs in $X_{0}$ are considered to be a fixed set of parameters. Clearly, (2.14) is equivalent to (1.12). We note that we are not introducing an activation function in the terminal layer.
Training this shallow network amounts to finding the optimizers $W_{j}^{**},b_{j}^{**}$ for the minimization problem
${\mathcal{C}}[W_{j}^{**},b_{j}^{**}]=\min_{W_{j},b_{j}}{\mathcal{C}}[W_{j},b_{j}]\,.$ (2.16)
Before we present our main results in the next section, we contrast our strategy with the commonly used approach via gradient descent.
2.1. Remarks on gradient descent
To determine the solution to the minimization problem (2.16),
${\mathcal{C}}[W_{j}^{**},b_{j}^{**}]=\min_{W_{j},b_{j}}\frac{1}{\sqrt{N}}\|W_{2}\sigma(W_{1}X_{0}+B_{1})+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}}\,,$ (2.17)
one aims to find the minimizers $W_{j}^{**},b_{j}^{**}$ that parametrize the optimally trained network. Our intention here is to explicitly construct (local) minimizers to the cost.
In contrast, minimization via the gradient descent algorithm is based on the dynamical system $({\mathcal{W}},\Phi_{\tau})$ where
${\mathcal{W}}:={\mathbb{R}}^{M^{2}}\times{\mathbb{R}}^{QM}\times{\mathbb{R}}^{M}\times{\mathbb{R}}^{Q}\;\;\;{\rm with}\;\;\underline{\theta}:=(\vec{W_{1}},\vec{W_{2}},b_{1},b_{2})\in{\mathcal{W}}$ (2.18)
is the space of weights and biases, and $\vec{A}:=(A_{1}^{T},\dots,A_{n}^{T})^{T}\in{\mathbb{R}}^{mn}$ with $A_{j}\in{\mathbb{R}}^{m}$ denoting the $j$-th column vector of the matrix $A\in{\mathbb{R}}^{m\times n}$. $\Phi_{\tau}$ denotes the flow generated by the gradient $G_{{\mathcal{C}}}[\underline{\theta}]:=-\nabla_{\underline{\theta}}{\mathcal{C}}[\underline{\theta}]$ of the cost function (or of its square), that is,
$\partial_{\tau}\underline{\theta}(\tau)=G_{{\mathcal{C}}}[\underline{\theta}(\tau)]\;\;\;,\;\;\underline{\theta}(\tau):=\Phi_{\tau}[\underline{\theta}_{0}]\;\;\;,\;\;\underline{\theta}_{0}\in{\mathcal{W}}\,,$ (2.19)
where $\tau\in{\mathbb{R}}_{+}$. Due to
$\partial_{\tau}{\mathcal{C}}[\underline{\theta}(\tau)]=G_{{\mathcal{C}}}[\underline{\theta}(\tau)]\cdot\partial_{t}\underline{\theta}(\tau)=-|G_{{\mathcal{C}}}[\underline{\theta}(\tau)]|^{2}\leq 0\,,$ (2.20)
the cost function ${\mathcal{C}}[\underline{\theta}(\tau)]$ decreases monotonically along orbits of $\Phi_{\tau}$. This is the basis of the majority of computational approaches to the determination of global minima or sufficiently good local minima. Gradient descent algorithms are highly sensitive to the choice of the initial data $\underline{\theta}_{0}$; the latter is in practice mostly picked at random, and vaguely informed by structural insights into the problem.
We distinguish the following situations for the approach via gradient descent.
  1. (1)
    If $W_{1},b_{1}$ are given suitable initial values so that along the gradient descent recursion, $W_{1}X_{0}+B_{1}$ have non-negative matrix components, then the activation function $\sigma$, simply acts as the identity ${\bf 1}_{M\times M}$. Accordingly, (2.17) can be determined by solving the least squares problem
    $\min_{W_{2},b_{2}}\|W_{2}X_{0}+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}}\,.$ (2.21)
    This is because we can redefine $W_{2}W_{1}\rightarrow W_{2}$ and $b_{2}\rightarrow b_{2}-W_{2}b_{1}$, as $W_{1}$ and $b_{1}$ are unknown. The solution is given by
    $W_{2}X_{0}=Y^{ext}{\mathcal{P}}\;\;\;,\;b_{2}=0$ (2.22)
    where ${\mathcal{P}}^{T}$ is the projector onto the range of $X_{0}^{T}$ (which has rank $Q$), and ${\mathcal{P}}^{\perp}={\bf 1}_{N\times N}-{\mathcal{P}}$.
  2. (2)
    The opposite extreme case is given when $W_{1}$ and $B_{1}$ are given initial values where all matrix components of $W_{1}X_{0}+B_{1}\in{\mathbb{R}}^{M\times N}$ are $\leq 0$, so that
    $\sigma(W_{1}X_{0}+B_{1})=0\,.$ (2.23)
    Then, the cost function becomes independent of $W_{1},b_{1}$, and its gradient in these variables is zero. One is left with the minimization problem
    $\min_{b_{2}}\|B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}}\,,$ (2.24)
    which has the solution
    $(b_{2})_{i}=\frac{1}{N}\sum_{j=1}^{N}[Y^{ext}]_{ij}=\sum_{j=1}^{Q}\frac{N_{j}}{N}y_{j}\,.$ (2.25)
    The corresponding trained network is not capable of matching inputs $x\in{\mathbb{R}}^{M}$ to outputs $y_{j}$.
  3. (3)
    An intermediate variant of the previous two cases is obtained when $W_{1}$ and $B_{1}$ attain values where some components $(b_{1})_{i_{\ell}}$, $\ell\in{\mathcal{L}}\subset\{1,\dots,M\}$, render the corresponding rows of $[W_{1}X_{0}+B_{1}]_{i_{\ell}j}\leq 0$ for $j=1,\dots,N$. Then, the cost function is independent of $(b_{1})_{i_{\ell}}$ and $[W_{1}]_{i_{\ell}j}$, $j=1,\dots,N$, $\ell\in{\mathcal{L}}$.
    Because by assumption, ${\rm rank}(Y^{ext})=Q$, we find that if
    ${\rm rank}(\sigma(W_{1}X_{0}+B_{1}))=Q\,,$ (2.26)
    the possibility of finding a solution to (2.17) remains intact, based on the surviving components of $W_{1},b_{1}$ that occupy the non-zero rows. In this paper, we will investigate aspects of this scenario.
  4. (4)
    On the other hand, in the same situation as in case (3), if
    ${\rm rank}(\sigma(W_{1}X_{0}+B_{1}))<Q\,,$ (2.27)
    then although one can minimize the cost function depending on the surviving variables, one will not find the global minimum (2.17).
  5. (5)
    Even if the initial values for $W_{1},b_{1}$ lead to case (1), the gradient descent flow might steer them into one of the scenarios in the cases (2) $\sim$ (4) where a rank loss is induced by $\sigma$. Once this occurs, the gradient flow becomes independent of the components of $W_{1},b_{1}$ with respect to which a rank loss is obtained; those components remain stationary (or plateau) at the values at which they are first eliminated by $\sigma$.
We note that a priori, only the bound
${\rm rank}(\sigma(W_{1}X_{0}+B_{1}))\leq{\rm rank}((W_{1}X_{0}+B_{1})\leq M\,,$ (2.28)
is available. In applications, one often finds $M>Q$ or $M\gg Q$. The scenario described in case (3) yields the maximal reduction of complexity to still ensure compatibility with the minimization problem (2.17).
The above discussion highlights the importance of the question whether $W_{1}$ and $b_{1}$ can be analytically determined in a manner that a maximal rank reduction from $M$ to $Q$ is accomplished, as in (2.26), while maintaining the solvability of (2.17). This is addressed in Theorem 3.1, below, based on the construction of accurate upper bounds on the cost function, without making any recourse to gradient flow algorithms.
3. Statement of Main Results
In this section, we present the main results of this paper, an explicit upper bound on the cost function in Theorem 3.1 which leads to the construction of an optimized constructively trained shallow network, and its natural geometric interpretation in Theorem 3.3. In Theorem 3.2, we construct an explicit degenerate local minimum of the cost function, for the special case of input and output space having equal dimension, $M=Q$. In Theorem 3.5, we address the effects of truncation of the training input matrix due to the activation function, in the case $Q=M$.
3.1. Upper bound on minimum of cost function for $M>Q$
Let
$\overline{x_{0,j}}:=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}x_{0,j,i}\in{\mathbb{R}}^{M}$ (3.1)
denote the average over the equivalence class of all training input vectors corresponding to the $j$-th output, and
$\Delta x_{0,j,i}:=x_{0,j,i}-\overline{x_{0,j}}\,.$ (3.2)
Moreover, let
$\overline{X_{0,j}}:=[\overline{x_{0,j}}\cdots\overline{x_{0,j}}]\in{\mathbb{R}}^{M\times N_{j}}$ (3.3)
and
$\overline{X_{0}}:=[\overline{X_{0,1}}\cdots\overline{X_{0,Q}}]\in{\mathbb{R}}^{M\times N}\,.$ (3.4)
Similarly, let
$\Delta X_{0,j}:=[\Delta x_{0,j,1}\cdots\Delta x_{0,j,N_{j}}]\in{\mathbb{R}}^{M\times N_{j}}\,.$ (3.5)
and
$\Delta X_{0}:=[\Delta X_{0,1}\cdots\Delta X_{0,Q}]\in{\mathbb{R}}^{M\times N}\,.$ (3.6)
Then,
$X_{0}=\overline{X_{0}}+\Delta X_{0}\,,$ (3.7)
and we let
$\delta:=\sup_{j,i}|\Delta x_{0,j,i}|\ \,.$ (3.8)
Next, we define the reduction of $\overline{X_{0}}$ as
$\overline{X_{0}^{red}}:=[\overline{x_{0,1}}\cdots\overline{x_{0,Q}}]\in{\mathbb{R}}^{M\times Q}$ (3.9)
and assume that its rank is maximal, ${\rm rank}(\overline{X_{0}^{red}})=Q$. The latter condition is necessary for the averages of training inputs to be able to distinguish between the different output values $y_{j}$, $j=1,\dots,Q$.
Because of ${\rm rank}(\overline{X_{0}^{red}})=Q$, the matrix
$(\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}}\in{\mathbb{R}}^{Q\times Q}$ (3.10)
is invertible, and we can define the Penrose inverse of $\overline{X_{0}^{red}}$,
${\rm Pen}[\overline{X_{0}^{red}}]:=\big{(}(\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}}\big{)}^{-1}(\overline{X_{0}^{red}})^{T}\in{\mathbb{R}}^{Q\times M}$ (3.11)
The orthoprojector onto the range of $\overline{X_{0}^{red}}$ is then given by
$P:=\overline{X_{0}^{red}}{\rm Pen}[\overline{X_{0}^{red}}]\in{\mathbb{R}}^{M\times M}$ (3.12)
and we note that ${\rm Pen}[\overline{X_{0}^{red}}]\overline{X_{0}^{red}}={\bf 1}_{Q\times Q}$. The projector property $P^{2}=P$ is thus easily verified, and orthogonality with respect to the Euclidean inner product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on ${\mathbb{R}}^{M}$ holds due to $P^{T}=P$, whereby $(v,Pw)=(Pv,w)$ for any $v,w\in{\mathbb{R}}^{M}$. In particular, we have
$P\overline{X_{0}^{red}}=\overline{X_{0}^{red}}\;\;\;{\rm and}\;\;\;P\overline{X_{0}}=\overline{X_{0}}\,,$ (3.13)
by construction.
Moreover, we define
$\delta_{P}:=\sup_{i,j}\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,j,i}\right|\,,$ (3.14)
which measures the relative size between $\overline{X_{0}^{red}}$ and $P\Delta X_{0}$, as $\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,j,i}\right|$ scales like the noise to signal ratio $\frac{|\Delta x|}{|x|}$ of training inputs.
Theorem 3.1.
Let $Q\leq M\leq QM<N$. Assume that $R\in O(M)$ diagonalizes $P,P^{\perp}$, and let $\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|$.
Let ${\mathcal{C}}[W_{i}^{*},b_{i}^{*}]$ be the cost function evaluated for the trained shallow network defined by the following weights and biases,
$W_{1}^{*}=R\,,$ (3.15)
and
$W_{2}^{*}=\widetilde{W}_{2}R^{T}\;\;\;{\rm where}\;\widetilde{W}_{2}=Y\;{\rm Pen}[\overline{X_{0}^{red}}]P\,.$ (3.16)
Moreover, $b_{1}^{*}=P_{R}b_{1}^{*}+P_{R}^{\perp}b_{1}^{*}$ with
$P_{R}b_{1}^{*}=\beta_{1}u_{M}\;\;,\;P_{R}^{\perp}b_{1}^{*}=-\delta P_{R}^{\perp}u_{M}\;\;\in{\mathbb{R}}^{M}\,,$ (3.17)
for $u_{M}\in{\mathbb{R}}^{M}$ as in (1.14), and
$b_{2}^{*}=-W_{2}^{*}P_{R}b_{1}^{*}\;\;\in{\mathbb{R}}^{Q}\,.$ (3.18)
Then, the minimum of the cost function satisfies the upper bound
$\min_{W_{j},b_{j}}{\mathcal{C}}[W_{j},b_{j}]\leq{\mathcal{C}}[W_{i}^{*},b_{i}^{*}]\leq\frac{1}{\sqrt{N}}\|Y\;{\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}\|_{{\mathcal{L}}^{2}}\,,$ (3.19)
which implies
$\min_{W_{j},b_{j}}{\mathcal{C}}[W_{j},b_{j}]\;\leq\;\|Y\|_{op}\;\delta_{P}\;$ (3.20)
where $N=\sum_{\ell=1}^{Q}N_{\ell}$.
We observe that, for any $\lambda>0$, the rescaling of training inputs
$x_{0,j,i}\rightarrow\lambda x_{0,j,i}\;\;\forall j=1,\dots,Q\;\;,\;i=1,\dots,N_{j}\,,$ (3.21)
induces
${\rm Pen}[\overline{X_{0}^{red}}]\rightarrow\lambda^{-1}{\rm Pen}[\overline{X_{0}^{red}}]\;\;\;,\;\Delta X_{0}\rightarrow\lambda\Delta X_{0}\,.$ (3.22)
Therefore, the upper bound in (3.19) is scaling invariant, as it only depends on the signal to noise ratio of training input data, which is controlled by $\delta_{P}$.
3.2. Exact degenerate local minimum in the case $M=Q$
We explicitly determine an exact local degenerate minimum of the cost function in the case $Q=M$ where the input and output spaces have the same dimension; here, we show that the upper bound obtained above differs from the sharp value by a relative error of order $O(\delta_{P}^{2})$. We note that in this situation, $P={\bf 1}_{Q\times Q}$ and $P^{\perp}=0$. Moreover, $\overline{X_{0}^{red}}$ is invertible, and therefore, ${\rm Pen}[\overline{X_{0}^{red}}]=(\overline{X_{0}^{red}})^{-1}$.
We prove this result relative to a weighted variant of the cost function defined as follows. Let ${\mathcal{N}}\in{\mathbb{R}}^{N\times N}$ be the block diagonal matrix given by
${\mathcal{N}}:={\rm diag}(N_{j}{\bf 1}_{N_{j}\times N_{j}}\,|\,j=1,\dots,Q)\,.$ (3.23)
We introduce the inner product on ${\mathbb{R}}^{Q\times N}$
$(A,B)_{{\mathcal{L}}^{2}_{\mathcal{N}}}:={\rm Tr}(A{\mathcal{N}}^{-1}B^{T})$ (3.24)
and
$\|A\|_{{\mathcal{L}}^{2}_{\mathcal{N}}}:=\sqrt{(A,A)_{{\mathcal{L}}^{2}_{\mathcal{N}}}}\,.$ (3.25)
We define the weighted cost function
${\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]:=\|X^{(2)}-Y^{ext}\|_{{\mathcal{L}}^{2}_{\mathcal{N}}}\,.$ (3.26)
This is equivalent to
${\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]:=\sqrt{\sum_{j=1}^{Q}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}|W_{2}\sigma(W_{1}x_{0,j,i}+b_{1})+b_{2}-y_{j}|^{2}_{{\mathbb{R}}^{Q}}}\,.$ (3.27)
The weights $\frac{1}{N_{j}}$ ensure that the contributions to the cost function belonging to inputs $y_{j}$ do not depend on their sample sizes $N_{j}$. We note that for uniform sample sizes, where $\frac{N_{j}}{N}=\frac{1}{Q}$ $\forall j$, we have that ${\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]=\sqrt{Q}{\mathcal{C}}[W_{i},b_{i}]$.
Theorem 3.2.
Assume $M=Q<MQ<N$. Then, if ${\rm rank}(\overline{X_{0}})=Q$, let
${\mathcal{P}}:={\mathcal{N}}^{-1}X_{0}^{T}(X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}\;\;\in{\mathbb{R}}^{N\times N}\,.$ (3.28)
Its transpose ${\mathcal{P}}^{T}$ is an orthonormal projector in ${\mathbb{R}}^{N}$ with respect to the inner product $\langle u,v\rangle_{{\mathcal{N}}}:=(u,{\mathcal{N}}^{-1}v)$. In particular, $(A{\mathcal{P}},B)_{{\mathcal{L}}^{2}_{\mathcal{N}}}=(A,B{\mathcal{P}})_{{\mathcal{L}}^{2}_{\mathcal{N}}}$ for all $A,B\in{\mathbb{R}}^{Q\times N}$. Letting ${\mathcal{P}}^{\perp}={\bf 1}_{N\times N}-{\mathcal{P}}$, the weighted cost function satisfies the upper bound
$\min_{W_{j},b_{j}}{\mathcal{C}}_{{\mathcal{N}}}[W_{j},b_{j}]$$\leq$ ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ (3.29)
$=$ $\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}$
$=$ $\big{\|}Y|\Delta_{2}^{rel}|^{\frac{1}{2}}\big{(}1+\Delta_{2}^{rel}\big{)}^{-\frac{1}{2}}\big{\|}_{{\mathcal{L}}^{2}}$
$\leq$ $(1-C_{0}\delta_{P}^{2})\;\|Y\;(\overline{X_{0}^{red}})^{-1}\Delta X_{0}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}\,,$
for a constant $C_{0}\geq 0$, and where
$\Delta_{2}^{rel}:=\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}\;\;,\;{\rm with}\;\Delta_{1}^{rel}:=(\overline{X_{0}^{red}})^{-1}\Delta X_{0}\,.$ (3.30)
The weights and biases realizing the upper bound are given by
$W_{1}^{*}={\bf 1}_{Q\times Q}$ (3.31)
and
$W_{2}^{*}=Y(\overline{X_{0}^{red}})^{T}(X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}\in{\mathbb{R}}^{Q\times Q}\,.$ (3.32)
Moreover,
$b_{1}^{*}=\beta_{1}u_{Q}\;\;,\;b_{2}^{*}=-W_{2}^{*}b_{1}^{*}\,,$ (3.33)
with $u_{Q}\in{\mathbb{R}}^{Q}$ as in (1.14), and where $\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|$ so that
$\sigma(W_{1}^{*}X_{0}+B_{1}^{*})=W_{1}^{*}X_{0}+B_{1}^{*}\,.$ (3.34)
Notably,
$W_{2}^{*}=\widetilde{W}_{2}+O(\delta_{P}^{2})$ (3.35)
differs by $O(\delta_{P}^{2})$ from $\widetilde{W}_{2}$ in (3.16) used for the upper bound (3.19).
In particular, ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is a local minimum of the cost function; it is degenerate, assuming the same value for all $W_{i},b_{i}$ such that $W_{1},b_{1}$ satisfy the condition (3.34).
Moreover, ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is invariant under reparametrizations of the training inputs $X_{0}\rightarrow KX_{0}$, for all $K\in GL(Q)$.
The proof is given in Section 5.
3.3. Geometric interpretation
To match an arbitrary test input $x\in{\mathbb{R}}^{M}$ with one of the output vectors $y_{j}$, $j\in\{1,\dots,Q\}$, let for $x\in{\mathbb{R}}^{M}$,
${\mathcal{C}}_{j}[x]:=|W_{2}^{*}(W_{1}^{*}x+b_{1}^{*})_{+}+b_{2}^{*}-y_{j}|\,.$ (3.36)
Here, $(\cdot)_{+}$ denotes the ramp function (2.9), acting component-wise on a vector.
Given an arbitrary test input $x\in{\mathbb{R}}^{M}$,
$j^{*}={\rm argmin}_{j}{\mathcal{C}}_{j}[x]$ (3.37)
implies that $x$ matches the $j^{*}$-th output $y_{j^{*}}$.
The proof of Theorem 3.1 provides us with a particular set of weights $W_{1}^{*},W_{2}^{*}$ and biases $b_{1}^{*},b_{2}^{*}$ through an explicit construction, which yield the upper bound (3.19). We will refer to the shallow network trained with this choice of $W_{i}^{*},b_{i}^{*}$ as the constructively trained shallow network. A detailed discussion is provided in Section 6.
Theorem 3.3.
Assume $Q\leq M\leq QM<N$. Let $W_{i}^{*},b_{i}^{*}$, $i=1,2$, denote the weights and biases determined in Theorem 3.1. Let
$\widetilde{W}_{2}:=Y\;{\rm Pen}[\overline{X_{0}^{red}}]\in{\mathbb{R}}^{Q\times M}\,,$ (3.38)
and define the metric
$d_{\widetilde{W}_{2}}(x,y):=|\widetilde{W}_{2}P(x-y)|\;\;\;{\rm for}\;x,y\in{\rm range}(P)$ (3.39)
on the $Q$-dimensional linear subspace ${\rm range}(P)\subset{\mathbb{R}}^{M}$, where ||$\cdot$|| ⋅ | denotes the Euclidean norm on ${\mathbb{R}}^{Q}$. Then,
${\mathcal{C}}_{j}[x]=d_{\widetilde{W}_{2}}(Px,\overline{x_{0,j}})\,,$ (3.40)
and matching an input $x\in{\mathbb{R}}^{M}$ with an output $y_{j^{*}}$ via the constructively trained shallow network is equivalent to the solution of the metric minimization problem
$j^{*}={\rm argmin}_{j\in\{1,\dots,Q\}}(d_{\widetilde{W}_{2}}(Px,\overline{x_{0,j}}))$ (3.41)
on the range of $P$.
The proof is given in Section 7. The geometric interpretation of the construction is as follows.
First of all, we note that $\widetilde{W}_{2}P=\widetilde{W}_{2}$ has full rank $Q$, therefore $x\mapsto|\widetilde{W}_{2}Px|^{2}$ is a non-degenerate quadratic form on the range of $P$. Therefore, (3.39) indeed defines a metric on the range of $P$.
The constructively trained shallow network thus obtained matches a test input $x\in{\mathbb{R}}^{M}$ with an output vector $y_{j^{*}}$ by splitting
$x=Px+P^{\perp}x\,,$ (3.42)
and by determining which of the average training input vectors $\overline{x_{0,j}}$, $j=1,\dots,Q$, is closest to $Px$ in the $d_{\widetilde{W}_{2}}$ metric. The trained shallow network cuts off the components $P^{\perp}x$.
3.4. Dependence on truncation
Our next result addresses the effect of truncations enacted by the activation function $\sigma$ in the case $M=Q$. For the results in Theorem 3.2, we assumed that $W_{1},b_{1}$ are in the region (depending on $X_{0}$) such that
$\sigma(W_{1}X_{0}+B_{1})=W_{1}X_{0}+B_{1}$ (3.43)
holds. Here we discuss the situation in which $\sigma$ acts nontrivially. In this case, we assume that $W_{1},b_{1}$ are given, with $W_{1}\in{\mathbb{R}}^{Q\times Q}$ invertible, and observe that all matrix components of
$X^{(1)}=\sigma(W_{1}X_{0}+B_{1})$ (3.44)
are non-negative, as $X^{(1)}$ lies in the image of $\sigma$, and hence,
$\sigma(X^{(1)})=X^{(1)}\,.$ (3.45)
We define the truncation map $\tau_{W_{1},b_{1}}$ as follows. It will play a major role in our explicit construction of local and global cost minimizers in Deep Learning networks in [4].
Definition 3.4.
Let $W_{1}\in GL(Q)$, $b_{1}\in{\mathbb{R}}^{Q}$. Then, the truncation map $\tau_{W_{1},b_{1}}:{\mathbb{R}}^{Q\times N}\rightarrow{\mathbb{R}}^{Q\times N}$ is defined by
$\tau_{W_{1},b_{1}}(X_{0})$ ::=:= $W_{1}^{-1}(\sigma(W_{1}X_{0}+B_{1})-B_{1})$ (3.46)
$=$ $W_{1}^{-1}(X^{(1)}-B_{1})\,.$
That is, $\tau_{W_{1},b_{1}}=a_{W_{1},b_{1}}^{-1}\circ\sigma\circ a_{W_{1},b_{1}}$ under the affine map $a_{W_{1},b_{1}}:X_{0}\mapsto W_{1}X_{0}+B_{1}$.
We say that $\tau_{W_{1},b_{1}}$ is rank preserving (with respect to $X_{0}$) if both
${\rm rank}(\tau_{W_{1},b_{1}}(X_{0}))$ $=$ ${\rm rank}(X_{0})$
${\rm rank}(\overline{\tau_{W_{1},b_{1}}(X_{0})})$ $=$ ${\rm rank}(\overline{X_{0}})$ (3.47)
hold, and that it is rank reducing otherwise.
Then, we verify that
$\sigma(W_{1}\tau_{W_{1},b_{1}}(X_{0})+B_{1})$ $=$ $\sigma(X^{(1)})=X^{(1)}$ (3.48)
$=$ $W_{1}\tau_{W_{1},b_{1}}(X_{0})+B_{1}\,.$
This means that while the condition (3.43) does not hold for $X_{0}$, it does hold for $\tau_{W_{1},b_{1}}(X_{0})$.
Accordingly, we can define the matrices
$\overline{\tau_{W_{1},b_{1}}(X_{0})}\;\;,\;\overline{(\tau_{W_{1},b_{1}}(X_{0}))^{red}}\;\;,\;\Delta(\tau_{W_{1},b_{1}}(X_{0}))$ (3.49)
in analogy to $\overline{X_{0}},\overline{X_{0}^{red}},\Delta X_{0}$.
If no rank reduction is induced by the truncation, $\overline{(\tau_{W_{1},b_{1}}(X_{0}))^{red}}$ is invertible, and we obtain the following theorem.
Theorem 3.5.
Let $Q=M$, and assume that the truncation map $\tau_{W_{1},b_{1}}(X_{0})$ is rank preserving. Then, for any fixed $(W_{1},b_{1})\in GL(Q)\times{\mathbb{R}}^{Q}$,
$\min_{W_{2},b_{2}}{\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]=\|Y^{ext}{\mathcal{P}}^{\perp}_{\tau_{W_{1},b_{1}}(X_{0})}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}$ (3.50)
where the projector ${\mathcal{P}}_{\tau_{W_{1},b_{1}}(X_{0})}$ is obtained from (3.28) by substituting the truncated training inputs $\tau_{W_{1},b_{1}}(X_{0})$ for $X_{0}$. In particular, one explicitly has
$\min_{W_{2},b_{2}}{\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]$ $=$ $\big{\|}Y|\Delta_{2}^{rel,tr}|^{\frac{1}{2}}\big{(}1+\Delta_{2}^{rel,tr}\big{)}^{-\frac{1}{2}}\big{\|}_{{\mathcal{L}}^{2}}$ (3.51)
$\leq$ $(1-C_{0}\delta_{P,tr}^{2})\|Y\Delta_{1}^{rel,tr}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}$
where
$\Delta_{2}^{rel,tr}$ ::=:= $\Delta_{1}^{rel,tr}{\mathcal{N}}^{-1}(\Delta_{1}^{rel,tr})^{T}\;\;,\;{\rm with}\;$
$\Delta_{1}^{rel,tr}$ ::=:= $(\overline{(\tau_{W_{1},b_{1}}(X_{0}))^{red}})^{-1}\Delta(\tau_{W_{1},b_{1}}(x_{0,j,i}))\,,$ (3.52)
in analogy to (3.29), and for some $C_{0}\geq 0$, where
$\delta_{P,tr}:=\sup_{j,i}\left|(\overline{(\tau_{W_{1},b_{1}}(X_{0}))^{red}})^{-1}\Delta(\tau_{W_{1},b_{1}}(x_{0,j,i}))\right|$ (3.53)
measures the signal to noise ratio of the truncated training input data.
Proof.
Due to (3.48), we may follow the proof of Theorem 3.2 verbatim, and we find that for the given fixed choice of $W_{1},b_{1}$, minimization in $W_{2},b_{2}$ yields (3.50) and (3.51). ∎
In general, (3.51) does not constitute a stationary solution with respect to $W_{1},b_{1}$, as the right hand sides of (3.50), (3.51) generically depend nontrivially on $W_{1},b_{1}$. It thus remains to determine the infimum of (3.50), (3.51) with respect to $W_{1},b_{1}$, in order to find a candidate for the global minimum. The latter, however, might depend sensitively on the detailed properties of $\Delta X_{0}$, which are random.
On the other hand, in the situation addressed in Theorem 3.2, $W_{1},b_{1}$ are contained in the parameter region ${\mathcal{F}}_{X_{0}}$ (depending on $X_{0}$) defined by the fixed point relation for the truncation map (equivalent to (3.43)),
${\mathcal{F}}_{X_{0}}:=\{(W_{1},b_{1})\in GL(Q)\times{\mathbb{R}}^{Q}\;|\;\tau_{W_{1},b_{1}}(X_{0})=X_{0}\}\,.$ (3.54)
As a result, we indeed obtained a degenerate stationary solution with respect to $W_{1},b_{1}$,
$\min_{(W_{1},b_{1})\in{\mathcal{F}}_{X_{0}};W_{2},b_{2}}{\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]=\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}$ (3.55)
in (3.29), with ${\mathcal{P}}$ given in (3.28). Hence, ${\mathcal{F}}_{X_{0}}$ parametrizes an invariant manifold of equilibria of the gradient descent flow, which corresponds to a level surface of the cost function.
Remark 3.6.
The global minimum of the cost function has either the form (3.50) for an optimizer $(W_{1}^{**},b_{1}^{**})$ (with a rank preserving truncation), or (3.55). This is because any other scenario employs a rank reducing truncation. Since $Y^{ext}$ has rank $Q$, minimization of (3.26) under the constraint ${\rm rank}(\tau_{W_{1},b_{1}}(X_{0}))<Q$ cannot yield a global minimum.
The key point in minimizing (3.50) with respect to $(W_{1},b_{1})$ is to determine a rank preserving truncation which minimizes the signal to noise ratio of the truncated training inputs. Whether for a local or global minimum of the form (3.50) or (3.55), respectively, a shallow network trained with the corresponding weights and biases $W_{i},b_{i}$, $i=1,2$, will be able to match inputs $x\in{\mathbb{R}}^{Q}$ to outputs $y_{j}$, through a straightforward generalization of Theorem 3.3.
We expect these considerations to carry over to the general case $Q<M$, but leave a detailed discussion to future work.
4. Proof of Theorem 3.1
Let $R\in O(M)$ diagonalize $P,P^{\perp}$ (which are symmetric),
$P_{R}:=RPR^{T}\;\;\;,\;\;\;P^{\perp}_{R}:=RP^{\perp}R^{T}\,,$ (4.1)
where $P_{R}$ has $Q$ diagonal entries of $1$, and all other components are 00, while $P^{\perp}_{R}$ has $M-Q$ diagonal entries of $1$, and all other components are 00. This does not characterize $R$ uniquely, as we may compose $R$ with an arbitrary $R^{\prime}\in O(M)$ that leaves the ranges of $P,P^{\perp}$ invariant. Diagonality of $P_{R},P_{R}^{\perp}$ implies that the activation function, which acts component-wise as the ramp function (2.9), satisfies
$\sigma(P_{R}v+P^{\perp}_{R}v)=\sigma(P_{R}v)+\sigma(P^{\perp}_{R}v)$ (4.2)
for any $v\in{\mathbb{R}}^{M}$, and where of course, $P_{R}+P_{R}^{\perp}={\bf 1}_{M\times M}$.
We write
$b_{1}=P_{R}b_{1}+P_{R}^{\perp}b_{1}$ (4.3)
where we choose
$P_{R}b_{1}=\beta_{1}u_{Q}\,,$ (4.4)
and we pick
$\beta_{1}\geq 2\rho\;\;\;,\;\;\rho:=\max_{j,i}|x_{0,j,i}|\,.$ (4.5)
This ensures that $RPx_{0,j,i}+P_{R}b_{1}\in B_{\rho}(\beta_{1}u_{Q})\subset P_{R}{\mathbb{R}}_{+}^{M}$ so that
$\sigma(RPx_{0,j,i}+P_{R}b_{1})=RPx_{0,j,i}+P_{R}b_{1}\;\;\;,\;\;\;j=1,\dots,Q\;\;,\;\forall\;i=1,\dots,N_{j}$ (4.6)
To construct the upper bound, we choose
$W_{1}=R\,.$ (4.7)
Then,
$X^{(1)}$ $=$ $\sigma(W_{1}X_{0}+B_{1})$ (4.8)
$=$ $\sigma((P_{R}+P_{R}^{\perp})RX_{0}+(P_{R}+P_{R}^{\perp})B_{1})$
$=$ $\sigma(P_{R}RX_{0}+P_{R}B_{1})+\sigma(P^{\perp}_{R}RX_{0}+P^{\perp}_{R}B_{1})$
$=$ $\sigma(RPX_{0}+P_{R}B_{1})+\sigma(RP^{\perp}\Delta X_{0}+P^{\perp}_{R}B_{1})\,.$
This is because by construction,
$P\overline{X_{0}}=\overline{X_{0}}\;\;\;,\;\;P^{\perp}\overline{X_{0}}=0\,.$ (4.9)
Moreover, we used
$P_{R}R=RP\;\;\;,\;\;P^{\perp}_{R}R=RP^{\perp}\,,$ (4.10)
from (4.1).
Going to the next layer, we have
$X^{(2)}=W_{2}X^{(1)}+B_{2}\,.$ (4.11)
Our goal is to minimize the expression on the right hand side of
$\sqrt{N}{\mathcal{C}}[W_{j},b_{j}]$$\leq$ $\|W_{2}\sigma(RPX_{0}+P_{R}B_{1})+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}}$ (4.12)
$\hskip 28.45274pt+\|W_{2}\sigma(RP^{\perp}\Delta X_{0}+P^{\perp}_{R}B_{1})\|_{{\mathcal{L}}^{2}}\,.$
We note that because ${\rm rank}(Y^{ext})=Q$, and
${\rm rank}(\sigma(\underbrace{RPX_{0}}_{=P_{R}RPX_{0}}+P_{R}B_{1}))\leq Q\,,$ (4.13)
minimization of the first term on the r.h.s. of (4.12) requires that the latter has maximal rank $Q$. Due to our choice of $R$ and $P_{R}b_{1}$ which imply that (4.6) holds, we find that
$\sigma(RPX_{0}+P_{R}B_{1})$ $=$ $RPX_{0}+P_{R}B_{1}$ (4.14)
$=$ $P_{R}R(\overline{X_{0}}+P\Delta X_{0})+P_{R}B_{1}\,.$
To minimize the first term on the r.h.s. of (4.12), one might want to try to solve
$W_{2}S=Y^{ext}$ (4.15)
where
$S:=P_{R}R(\overline{X_{0}}+P\Delta X_{0})+P_{R}B_{1}\in{\mathbb{R}}^{M\times N}$ (4.16)
has rank $Q<M$ (due to $P_{R}$). However, this implies that the $QM$ unknown matrix components of $W_{2}\in{\mathbb{R}}^{Q\times M}$ are determined by $QN>QM$ equations. This is an overdetermined problem which generically has no solution. However, we can solve the problem to zeroth order in $\Delta X_{0}$,
$W_{2}S_{0}=Y^{ext}$ (4.17)
with
$S_{0}:=S|_{\Delta X_{0}=0}=P_{R}R\overline{X_{0}}=RP\overline{X_{0}}$ (4.18)
(recalling from (3.4) that $\overline{X_{0}}$ only contains $Q$ distinct column vectors $\overline{x_{0,j}}$ and their identical copies) and subsequently minimize the resulting expression in the first term on the r.h.s. of (4.12) with respect to $P_{R}B_{1}$.
This is equivalent to requiring that $W_{2}\in{\mathbb{R}}^{Q\times M}$ solves
$W_{2}RP\overline{X_{0}^{red}}=Y\,.$ (4.19)
To this end, we make the ansatz
$W_{2}=A(R\overline{X_{0}^{red}})^{T}$ (4.20)
for some $A\in{\mathbb{R}}^{Q\times Q}$. Then, recalling that $P\overline{X_{0}^{red}}=\overline{X_{0}^{red}}$,
$A(R\overline{X_{0}^{red}})^{T}(R\overline{X_{0}^{red}})=A(\overline{X_{0}^{red}})^{T}(\overline{X_{0}^{red}})=Y\,,$ (4.21)
so that solving for $A$, we get
$W_{2}$ $=$ $Y((\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}})^{-1}(R\overline{X_{0}^{red}})^{T}$ (4.22)
$=$ $\widetilde{W}_{2}R^{T}\,$
with $\widetilde{W}_{2}$ as defined in (3.38). Hereby, $W_{2}$ is fully determined by $Y$ and $\overline{X_{0}^{red}}$ (as $\overline{X_{0}^{red}}$ determines $P$ and $R$).
Next, we can choose $P^{\perp}_{R}b_{1}$ in a suitable manner that
$\sigma(RP^{\perp}\Delta X_{0}+P^{\perp}_{R}B_{1})=0\,.$ (4.23)
This can be accomplished with
$P^{\perp}_{R}b_{1}=-\delta P^{\perp}_{R}u_{M}\in{\mathbb{R}}^{M}$ (4.24)
where $\delta$ is defined in (3.8), and
$u_{M}:=(1,1,\dots,1)^{T}\in{\mathbb{R}}^{M}\,.$ (4.25)
With
$E:=u_{M}u_{N}^{T}=[u_{M}\cdots u_{M}]\in{\mathbb{R}}^{M\times N}\,,$ (4.26)
this ensures that the argument on the l.h.s. of (4.23), given by
$RP^{\perp}\Delta X_{0}+P^{\perp}_{R}B_{1}=P^{\perp}_{R}R\Delta X_{0}-\delta P^{\perp}_{R}E\,,$ (4.27)
is a matrix all of whose components are $\leq 0$. Therefore, $\sigma$ maps it to 00.
Consequently, we find
$\sqrt{N}{\mathcal{C}}[W_{j},b_{j}]\leq\|W_{2}(RP\Delta X_{0}+P_{R}B_{1})+B_{2}\|_{{\mathcal{L}}^{2}}\,.$ (4.28)
To minimize the r.h.s., we may bias
$b_{2}\rightarrow b_{2}-W_{2}P_{R}b_{1}$ (4.29)
so that it remains to minimize the r.h.s. of
$\sqrt{N}{\mathcal{C}}[W_{j},b_{j}]\leq\|W_{2}RP\Delta X_{0}+B_{2}\|_{{\mathcal{L}}^{2}}\,.$ (4.30)
with respect to
$B_{2}=[b_{2}\cdots b_{2}]=b_{2}u_{N}^{T}\;\;\in{\mathbb{R}}^{Q\times N}\,.$ (4.31)
We find
00 0 $=$ $\partial_{b_{2}}\|W_{2}RP\Delta X_{0}+\beta_{2}u_{N}^{T}\|_{{\mathcal{L}}^{2}}^{2}$ (4.32)
$=$ $\partial_{b_{2}}\Big{(}2{\rm Tr}(W_{2}RP\underbrace{\Delta X_{0}u_{N}}_{=0}b_{2}^{T})+{\rm Tr}(u_{N}\underbrace{b_{2}^{T}b_{2}}_{=|b_{2}|^{2}}u_{N}^{T})\Big{)}\,,$
using cyclicity of the trace. The first term inside the bracket on the second line is zero because for every $i\in\{1,\dots,M\}$,
$(\Delta X_{0}u_{N})_{i}$ $=$ $\sum_{\ell=1}^{N}[\Delta X_{0}]_{i\ell}$ (4.33)
$=$ $\sum_{j=1}^{Q}\sum_{\ell=1}^{N_{j}}(x_{0,j,\ell}-\overline{x_{0,j}})_{i}$
$=$ $0\,.$
The second term inside the bracket in (4.32) is proportional to $|b_{2}|^{2}$, so that its gradient is proportional to $b_{2}$, and vanishes for $b_{2}=0$. Hence, $B_{2}=0$, which means that
$B_{2}=-W_{2}P_{R}B_{1}$ (4.34)
for the original variable $b_{2}$ prior to the bias (5.10).
We thus arrive at an upper bound
$\min_{W_{j},b_{j}}\sqrt{N}{\mathcal{C}}[W_{j},b_{j}]$$\leq$ $\|\widetilde{W}_{2}P\Delta X_{0}\|_{{\mathcal{L}}^{2}}$ (4.35)
$=$ $\|Y{\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}\|_{{\mathcal{L}}^{2}}\,.$
This establishes the asserted upper bound in (3.19).
Next, we bound
$\|\widetilde{W}_{2}P\Delta X_{0}\|_{{\mathcal{L}}^{2}}$$\leq$ $\|Y\|_{op}\|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}\|_{{\mathcal{L}}^{2}}\,.$ (4.36)
Let for notational brevity $A:={\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}$. We have
$\|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}\|_{{\mathcal{L}}^{2}}^{2}$ $=$ ${\rm Tr}(AA^{T})$ (4.37)
$=$ ${\rm Tr}(A^{T}A)$
$=$ $\sum_{\ell=1}^{Q}\sum_{i=1}^{N_{j}}|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,\ell,i}|^{2}$
$\leq$ $\Big{(}\sup_{i,\ell}\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,\ell,i}\right|^{2}\Big{)}\sum_{\ell=1}^{Q}\sum_{i=1}^{N_{\ell}}1$
$=$ $N\sup_{i,\ell}\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,\ell,i}\right|^{2}\,.$
Therefore,
$\min_{W_{j},b_{j}}\sqrt{N}{\mathcal{C}}[W_{j},b_{j}]$$\leq$ $\|\widetilde{W}_{2}P\Delta X_{0}\|_{{\mathcal{L}}^{2}}$ (4.38)
$\leq$ $\sqrt{N}\;\|Y\|_{op}\sup_{i,\ell}\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,\ell,i}\right|$
$\leq$ $\sqrt{N}\;\|Y\|_{op}\delta_{P}\,,$
which proves (3.20). This concludes the proof of Theorem 3.1. ∎
5. Proof of Theorem 3.2
To prove the exact degenerate local minimum in the special case $M=Q$, we note that in this situation,
$P={\bf 1}_{Q\times Q}\;\;\;,\;P^{\perp}=0\,.$ (5.1)
In place of (4.12), the cost function is given by
${\mathcal{C}}_{{\mathcal{N}}}[W_{j},b_{j}]$ $=$ $\|W_{2}\sigma(W_{1}(\overline{X_{0}}+\Delta X_{0})+B_{1})+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}_{\mathcal{N}}}\,,$ (5.2)
defined in (3.26), which is weighted by the inverse of the block diagonal matrix
${\mathcal{N}}={\rm diag}(N_{j}{\bf 1}_{N_{j}\times N_{j}}\,|\,j=1,\dots,Q)\;\in{\mathbb{R}}^{N\times N}\,.$ (5.3)
Using a similar choice for $B_{1}$ as in (4.4), we may assume that the matrix components of $W_{1}(\overline{X_{0}}+\Delta X_{0})+B_{1}$ are non-negative. To be precise, we let
$b_{1}=\beta_{1}u_{Q}$ (5.4)
for some $\beta_{1}\geq 0$ large enough that
$x_{0,j,i}+\beta_{1}u_{Q}\in{\mathbb{R}}_{+}^{Q}$ (5.5)
for all $j=1,\dots,Q$, $i=1,\dots,N_{j}$. This is achieved with
$\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|\,.$ (5.6)
Hence, choosing
$W_{1}={\bf 1}_{Q\times Q}\,,$ (5.7)
we find
$\sigma(W_{1}X_{0}+B_{1})=\overline{X_{0}}+\Delta X_{0}+B_{1}\,.$ (5.8)
Thus, 5.2 implies
$\min_{W_{j},b_{j}}{\mathcal{C}}_{{\mathcal{N}}}[W_{j},b_{j}]\leq\|W_{2}(\overline{X_{0}}+\Delta X_{0})+W_{2}B_{1}+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}_{\mathcal{N}}}\,.$ (5.9)
To minimize the r.h.s., we note that since $W_{2}$ and $B_{2}$ are both unknown, we can redefine $W_{2}\rightarrow W_{2}$, and we can bias
$b_{2}\rightarrow b_{2}-W_{2}b_{1}\,.$ (5.10)
Therefore, we find
$\min_{W_{j},b_{j}}{\mathcal{C}}_{{\mathcal{N}}}[W_{j},b_{j}]\leq\min_{W_{2},b_{2}}\|W_{2}(\overline{X_{0}}+\Delta X_{0})+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}_{\mathcal{N}}}\,.$ (5.11)
Minimizing with respect to $W_{2}\in{\mathbb{R}}^{Q\times Q}$ is a least squares problem whose solution has the form
$(W_{2}(\overline{X_{0}}+\Delta X_{0})-Y^{ext}_{b_{2}}){\mathcal{P}}=0\,,$ (5.12)
where
$Y^{ext}_{b_{2}}:=Y^{ext}-B_{2}\,,$ (5.13)
${\mathcal{P}}^{T}$ is a projector onto the range of $X_{0}^{T}$, and remains to be determined. To solve (5.12), we require $W_{2}$ to satisfy
$W_{2}(\overline{X_{0}}+\Delta X_{0}){\mathcal{N}}^{-1}(\overline{X_{0}}+\Delta X_{0})^{T}$ (5.14)
$\hskip 56.9055pt=Y^{ext}_{b_{2}}{\mathcal{N}}^{-1}(\overline{X_{0}}+\Delta X_{0})^{T}\;\;\in{\mathbb{R}}^{Q\times Q}\,.$
Because
$\overline{X_{0}}$ $=$ $[\overline{x_{0,1}}u_{N_{1}}^{T}\cdots\overline{x_{0,Q}}u_{N_{Q}}^{T}]\,,$
$Y^{ext}_{b_{2}}$ $=$ $[(y_{1}-b_{2})u_{N_{1}}^{T}\cdots(y_{Q}-b_{2})u_{N_{Q}}^{T}]\,,$ (5.15)
where
$u_{N_{j}}:=(1,1,\dots,1)^{T}\;\;\in{\mathbb{R}}^{N_{j}}\,,$ (5.16)
we find that
$\Delta X_{0}{\mathcal{N}}^{-1}(\overline{X_{0}})^{T}$ $=$ $\sum_{j=1}^{Q}\frac{1}{N_{j}}\Delta X_{0,j}u_{N_{j}}\overline{x_{0,j}}^{T}$ (5.17)
$=$ $\sum_{j=1}^{Q}\underbrace{\Big{(}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\Delta x_{0,j,i}\Big{)}}_{=0}\overline{x_{0,j}}^{T}$
$=$ $0\,.$
For the same reason,
$\overline{X_{0}}{\mathcal{N}}^{-1}\Delta X_{0}^{T}=0\;\;\;,\;Y^{ext}{\mathcal{N}}^{-1}\Delta X_{0}^{T}=0\,.$ (5.18)
Moreover,
$Y^{ext}_{b_{2}}{\mathcal{N}}^{-1}\overline{X_{0}}^{T}$ $=$ $\sum_{j=1}^{Q}\frac{1}{N_{j}}(y_{j}-b_{2})\underbrace{u_{N_{j}}^{T}u_{N_{j}}}_{=N_{j}}\overline{x_{0,j}}^{T}$ (5.19)
$=$ $Y_{b_{2}}\overline{X_{0}^{red}}^{T}\,,$
where
$Y_{b_{2}}=[(y_{1}-b_{2})\cdots(y_{Q}-b_{2})]\;\;\in{\mathbb{R}}^{Q\times Q}$ (5.20)
and
$\overline{X_{0}}{\mathcal{N}}^{-1}\overline{X_{0}}^{T}=\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}\,.$ (5.21)
Therefore, (5.14) reduces to
$W_{2}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}+\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})=Y_{b_{2}}\overline{X_{0}^{red}}^{T}\,,$ (5.22)
and invertibility of $\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}\in{\mathbb{R}}^{Q\times Q}$ implies invertibility of the matrix in brackets on the l.h.s. of (5.22), because
$D_{2}[X_{0}]:=\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T}=\sum_{j=1}^{Q}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\Delta x_{0,j,i}\Delta x_{0,j,i}^{T}\;\;\in{\mathbb{R}}^{Q\times Q}\,$ (5.23)
is a non-negative operator. Hence,
$W_{2}=Y_{b_{2}}\overline{X_{0}^{red}}^{T}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}+\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})^{-1}\,,$ (5.24)
so that
$W_{2}X_{0}=Y^{ext}_{b_{2}}{\mathcal{P}}\,.$ (5.25)
Here,
${\mathcal{P}}:={\mathcal{N}}^{-1}X_{0}^{T}(X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}\in{\mathbb{R}}^{N\times N}$ (5.26)
is a rank $Q$ projector ${\mathcal{P}}^{2}={\mathcal{P}}$ which is orthogonal with respect to the inner product on ${\mathbb{R}}^{N}$ defined by ${\mathcal{N}}$, i.e., ${\mathcal{P}}^{T}{\mathcal{N}}={\mathcal{N}}{\mathcal{P}}$. ${\mathcal{P}}^{T}$ is the projector onto the range of $X_{0}^{T}$. We note that with respect to the inner product defined with ${\mathcal{N}}^{-1}$ (which we are using here), we have ${\mathcal{N}}^{-1}{\mathcal{P}}^{T}={\mathcal{P}}{\mathcal{N}}^{-1}$.
Next, in order to control $W_{2}X_{0}-Y^{ext}_{b_{2}}=-Y^{ext}_{b_{2}}{\mathcal{P}}^{\perp}$, we observe that
$Y^{ext}_{b_{2}}{\mathcal{P}}$ $=$ $Y^{ext}_{b_{2}}{\mathcal{N}}^{-1}X_{0}^{T}(X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}$
$=$ $Y_{b_{2}}\overline{X_{0}^{red}}^{T}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}+D_{2}[X_{0}])^{-1}X_{0}$
$=$ $Y_{b_{2}}\overline{X_{0}^{red}}^{T}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T})^{-1}X_{0}$
$-Y_{b_{2}}\overline{X_{0}^{red}}^{T}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T})^{-1}D_{2}[X_{0}](X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}$
$=$ $Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}X_{0}$
$-Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}D_{2}[X_{0}](X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}$
where we have used the matrix identity
$(A+B)^{-1}$ $=$ $A^{-1}-A^{-1}B(A+B)^{-1}$ (5.28)
$=$ $A^{-1}-(A+B)^{-1}BA^{-1}$
for $A$, $A+B$ invertible. We observe that
$Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}X_{0}$ $=$ $Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}\overline{X_{0}}+Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}\Delta{X_{0}}$ (5.29)
$=$ $Y^{ext}_{b_{2}}+Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}\Delta{X_{0}}\,.$
To pass to the last line, we used that $(\overline{X_{0}^{red}})^{-1}\overline{X_{0}}=\,[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]$ where $\{e_{j}\in{\mathbb{R}}^{Q}\,|\,j=1,\dots,Q\}$ are the unit basis vectors.
We therefore conclude that, due to ${\mathcal{P}}^{\perp}={\bf 1}-{\mathcal{P}}$,
$Y^{ext}_{b_{2}}{\mathcal{P}}^{\perp}$ $=$ $-Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}\Delta{X_{0}}+{\mathcal{R}}$ (5.30)
where
${\mathcal{R}}$ ::=:= $Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}D_{2}[X_{0}](X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}$ (5.31)
satisfies
${\mathcal{R}}{\mathcal{P}}={\mathcal{R}}\,,$ (5.32)
as can be easily verified.
At this point, we have found that
$\min_{W_{j},b_{j}}{\mathcal{C}}_{{\mathcal{N}}}[W_{j},b_{j}]\leq\min_{b_{2}}\|Y^{ext}_{b_{2}}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}\,.$ (5.33)
To minimize with respect to $b_{2}$, we set
00 0 $=$ $\partial_{b_{2}}\|Y^{ext}_{b_{2}}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}$ (5.34)
$=$ $\partial_{b_{2}}\Big{(}-2{\rm Tr}\Big{(}Y^{ext}{\mathcal{P}}^{\perp}{\mathcal{N}}^{-1}({\mathcal{P}}^{\perp})^{T}u_{N}b_{2}^{T}\Big{)}+{\rm Tr}\big{(}b_{2}u_{N}^{T}{\mathcal{P}}^{\perp}{\mathcal{N}}^{-1}({\mathcal{P}}^{\perp})^{T}u_{N}b_{2}^{T}\big{)}\Big{)}$
$=$ $\partial_{b_{2}}\Big{(}-2{\rm Tr}\Big{(}b_{2}^{T}Y^{ext}{\mathcal{P}}^{\perp}{\mathcal{N}}^{-1}({\mathcal{P}}^{\perp})^{T}u_{N}\Big{)}+|{\mathcal{N}}^{-1/2}({\mathcal{P}}^{\perp})^{T}u_{N}|^{2}|b_{2}|^{2}\Big{)}$
$=$ $\partial_{b_{2}}\Big{(}-2{\rm Tr}\Big{(}b_{2}^{T}Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}\underbrace{\Delta{X_{0}}{\mathcal{N}}^{-1}u_{N}}_{=0}\Big{)}\;+\;{\rm Tr}\Big{(}b_{2}^{T}\underbrace{{\mathcal{R}}{\mathcal{N}}^{-1}({\mathcal{P}}^{\perp})^{T}}_{={\mathcal{R}}{\mathcal{P}}{\mathcal{N}}^{-1}{\mathcal{P}}^{\perp}=0}u_{N}\Big{)}$
$\hskip 56.9055pt+|{\mathcal{N}}^{-1/2}({\mathcal{P}}^{\perp})^{T}u_{N}|^{2}|b_{2}|^{2}\Big{)}$
using ${\mathcal{R}}={\mathcal{R}}{\mathcal{P}}$ from (5.32) so that ${\mathcal{R}}{\mathcal{N}}^{-1}{\mathcal{P}}^{\perp}={\mathcal{R}}{\mathcal{P}}{\mathcal{N}}^{-1}{\mathcal{P}}^{\perp}={\mathcal{R}}{\mathcal{P}}{\mathcal{P}}^{\perp}{\mathcal{N}}^{-1}=0$, and recalling $u_{N}=(1,1,\dots,1)^{T}\in{\mathbb{R}}^{N}$. This implies that
$b_{2}=0\,.$ (5.35)
We note that this corresponds to
$b_{2}=-W_{2}b_{1}$ (5.36)
prior to the bias (5.10), and hence arrive at (3.33), as claimed.
Next, we derive a simplified expression for $\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}$. For notational convenience, let
$\Delta_{1}^{rel}$ ::=:= $(\overline{X_{0}^{red}})^{-1}\Delta X_{0}$
$\Delta_{2}^{rel}$ ::=:= $(\overline{X_{0}^{red}})^{-1}D_{2}[X_{0}](\overline{X_{0}^{red}})^{-T}$ (5.37)
$=$ $\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}$
where we recall (5.23). Then, we obtain (with $b_{2}=0$)
${\mathcal{R}}$ $=$ $Y\Delta_{2}^{rel}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}(\overline{X_{0}^{red}})^{-1}X_{0}$ (5.38)
$=$ $Y\Delta_{2}^{rel}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\Big{(}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]+\Delta_{1}^{rel}\Big{)}$
where we used $X_{0}{\mathcal{N}}^{-1}X_{0}^{T}=\overline{X_{0}^{red}}(1+\Delta_{2}^{rel})(\overline{X_{0}^{red}})^{T}$ in (5.31) to obtain the first line. We find that
$-Y^{ext}{\mathcal{P}}^{\perp}$ (5.39)
$=$ $Y\Delta_{1}^{rel}-Y\Delta_{2}^{rel}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\Big{(}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]+\Delta_{1}^{rel}\Big{)}$
$=$ $Y(1-\Delta_{2}^{rel}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1})\Delta_{1}^{rel}-Y\Delta_{2}^{rel}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]$
$=$ $Y(1+\Delta_{2}^{rel})^{-1}\Delta_{1}^{rel}-Y\Delta_{2}^{rel}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]\,.$
Therefore, we obtain that
$\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}$ $=$ $(I)+(II)+(III)$ (5.40)
where
$(I)$ ::=:= ${\rm Tr}\Big{(}Y(1+\Delta_{2}^{rel})^{-1}\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}(1+\Delta_{2}^{rel})^{-T}Y^{T}\Big{)}$ (5.41)
$=$ ${\rm Tr}\Big{(}Y(1+\Delta_{2}^{rel})^{-1}\Delta_{2}^{rel}(1+\Delta_{2}^{rel})^{-1}Y^{T}\Big{)}$
where we note that $\Delta_{2}^{rel}=(\Delta_{2}^{rel})^{T}\in{\mathbb{R}}^{Q\times Q}$ is symmetric. Moreover,
$(II)$ ::=:= $-2{\rm Tr}\Big{(}Y(1+\Delta_{2}^{rel})^{-1}\Delta_{1}^{rel}{\mathcal{N}}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]^{T}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\Delta_{2}^{rel}Y^{T}\Big{)}$ (5.42)
$=$ $-\sum_{j=1}^{Q}2{\rm Tr}\Big{(}Y(1+\Delta_{2}^{rel})^{-1}(\overline{X_{0}^{red}})^{-1}\underbrace{\frac{1}{N_{j}}\Delta X_{0,j}u_{N_{j}}e_{j}^{T}}_{=0}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\Delta_{2}^{rel}Y^{T}\Big{)}$
$=$ 00 0
where we used that $\frac{1}{N_{j}}\Delta X_{0,j}u_{N_{j}}=\overline{\Delta X_{0,j}}=0$ for all $j=1,\dots,Q$. Furthermore,
$(III)$ ::=:= ${\rm Tr}\Big{(}Y\Delta_{2}^{rel}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]{\mathcal{N}}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]^{T}$ (5.43)
$\hskip 56.9055pt\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\Delta_{2}^{rel}Y^{T}\Big{)}$
$=$ ${\rm Tr}\Big{(}Y\Delta_{2}^{rel}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\sum_{j=1}^{Q}\underbrace{\frac{1}{N_{j}}u_{N_{j}}^{T}u_{N_{j}}}_{=1}e_{j}e_{j}^{T}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\Delta_{2}^{rel}Y^{T}\Big{)}$
$=$ ${\rm Tr}\Big{(}Y\Delta_{2}^{rel}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\Delta_{2}^{rel}Y^{T}\Big{)}$
$=$ ${\rm Tr}\Big{(}Y\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}(\Delta_{2}^{rel})^{2}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}Y^{T}\Big{)}$
using $\sum_{j=1}^{Q}e_{j}e_{j}^{T}={\bf 1}_{Q\times Q}$, and commutativity of $\Delta_{2}^{rel}$ and $(1+\Delta_{2}^{rel})^{-1}$.
We conclude that
$\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}$ $=$ $(I)+(III)$ (5.44)
$=$ ${\rm Tr}\Big{(}Y\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\Delta_{2}^{rel}(1+\Delta_{2}^{rel})\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}Y^{T}\Big{)}$
$=$ ${\rm Tr}\Big{(}Y\Delta_{2}^{rel}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}Y^{T}\Big{)}$
$=$ ${\rm Tr}\Big{(}Y|\Delta_{2}^{rel}|^{\frac{1}{2}}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}|\Delta_{2}^{rel}|^{\frac{1}{2}}Y^{T}\Big{)}$
$\leq$ $\|\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\|_{op}\;{\rm Tr}\Big{(}Y\Delta_{2}^{rel}Y^{T}\Big{)}$
since the matrix $\Delta_{2}^{rel}\geq 0$ is positive semidefinite. We note that
${\rm Tr}\Big{(}Y\Delta_{2}^{rel}Y^{T}\Big{)}$ $=$ ${\rm Tr}\Big{(}Y\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}Y^{T}\Big{)}$ (5.45)
$=$ $\|Y(\overline{X_{0}^{red}})^{-1}\Delta{X_{0}}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}\,.$
Moreover, letting $\lambda_{-}:=\inf{\rm spec}(\Delta_{2}^{rel})$ and $\lambda_{+}:=\sup{\rm spec}(\Delta_{2}^{rel})$ denote the smallest and largest eigenvalue of the non-negative, symmetric matrix $\Delta_{2}^{rel}\geq 0$,
$0\leq\lambda_{-}\leq\lambda_{+}=\|\Delta_{2}^{rel}\|_{op}=\|\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}\|_{op}\leq C\delta_{P}^{2}\,.$ (5.46)
Therefore,
$\|\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\|_{op}=(1+\lambda_{-})^{-1}\leq 1-C_{0}^{\prime}\delta_{P}^{2}$ (5.47)
for a constant $C_{0}^{\prime}=0$ if ${\rm rank}(\Delta_{2}^{rel})<Q$ whereby $\lambda_{-}=0$, and $C_{0}^{\prime}>0$ if ${\rm rank}(\Delta_{2}^{rel})=Q$.
We find the following bound from (5.44),
$\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}$$\leq$ $\sqrt{\|\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\|_{op}\;{\rm Tr}\Big{(}Y\Delta_{2}^{rel}Y^{T}\Big{)}}$ (5.48)
$\leq$ $(1-C_{0}\delta_{P}^{2})\;\|Y(\overline{X_{0}^{red}})^{-1}\Delta{X_{0}}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}\,,$
as claimed, where $C_{0}$ is proportional to $C_{0}^{\prime}$.
Finally, we use that
$\widetilde{W}_{2}$ $=$ $Y\overline{X_{0}^{red}}^{T}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T})^{-1}$ (5.49)
$=$ $Y(\overline{X_{0}^{red}})^{-1}\,$
agrees with (3.38) because $\overline{X_{0}^{red}}$ is invertible, and hence $(\overline{X_{0}^{red}})^{-1}={\rm Pen}[\overline{X_{0}^{red}}]$. Hence, we observe that (5.24) implies
$W_{2}$ $=$ $\widetilde{W}_{2}-\widetilde{W}_{2}(\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}+\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})^{-1}$ (5.50)
$=$ $\widetilde{W}_{2}-Y(\overline{X_{0}^{red}})^{-1}(\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})(\overline{X_{0}^{red}})^{-T}$
$\hskip 56.9055pt\Big{(}{\bf 1}+(\overline{X_{0}^{red}})^{-1}(\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})(\overline{X_{0}^{red}})^{-T}\Big{)}^{-1}(\overline{X_{0}^{red}})^{-1}$
$=$ $\widetilde{W}_{2}-Y\Delta_{2}^{rel}({\bf 1}+\Delta_{2}^{rel})^{-1}(\overline{X_{0}^{red}})^{-1}$
Therefore, in operator norm,
$\|W_{2}-\widetilde{W}_{2}\|_{op}$$\leq$ $\|Y\|_{op}\|\Delta_{2}^{rel}\|_{op}\|({\bf 1}+\Delta_{2}^{rel})^{-1}\|_{op}\|(\overline{X_{0}^{red}})^{-1}\|_{op}$ (5.51)
$\leq$ $C\delta_{P}^{2}$
where we used $\|Y\|_{op}<C$, $\|\Delta_{2}^{rel}\|_{op}<C\delta_{P}^{2}$, $\|({\bf 1}+\Delta_{2}^{rel})^{-1}\|_{op}\leq 1$, $\|(\overline{X_{0}^{red}})^{-1}\|_{op}<C$ to obtain the last line.
We note that when $M>Q$, it follows that $\overline{X_{0}^{red}}(\overline{X_{0}^{red}})^{T}$ is an $M\times M$ matrix of rank $Q$, and is hence not invertible, so that the above arguments do not apply.
From here on, we denote the weights and biases determined above in (5.7), (5.24) and in (5.4), (5.36), by $W_{1}^{*},W_{2}^{*}$ and $b_{1}^{*},b_{2}^{*}$, respectively. To prove that ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is a local minimum of the cost function ${\mathcal{C}}_{{\mathcal{N}}}$, we observe that for
$\rho:=\max_{j,i}|x_{0,j,i}|$ (5.52)
and
$\beta_{1}\geq 2\rho\;\;\;,\;\;b_{1}^{*}=\beta_{1}u_{Q}\,,$ (5.53)
we have
$W_{1}^{*}x_{0,j,i}+b_{1}^{*}$ $=$ $x_{0,j,i}+b_{1}^{*}$ (5.54)
$\in$ $B_{\rho}(\beta_{1}u_{Q})\subset{\mathbb{R}}_{+}^{Q}\,.$
$\forall j=1,\dots,Q$ and $\forall\;i=1,\dots,N_{j}$. That is, all training input vectors are contained in a ball of radius $\rho$ centered at a point on the diagonal $\beta_{1}u_{Q}\in{\mathbb{R}}^{Q}$ with $\beta_{1}\geq 2\rho$. This ensures that the coordinates of all $W_{1}^{*}x_{0,j,i}+b_{1}^{*}$ are strictly positive, and therefore,
$\sigma(W_{1}^{*}X_{0}+B_{1}^{*})=W_{1}^{*}X_{0}+B_{1}^{*}\,.$ (5.55)
For $\epsilon>0$ small, we consider an infinitesimal translation
$b_{1}^{*}\rightarrow b_{1}^{*}+\widetilde{b}_{1}$ (5.56)
combined with an infinitesimal transformation
$W_{1}^{*}\rightarrow W_{1}^{*}\widetilde{W}_{1}$ (5.57)
where $|\widetilde{b}_{1}|<\epsilon$ and $\|\widetilde{W}_{1}-{\bf 1}_{Q\times Q}\|_{op}<\epsilon$. Then, we have that
$W_{1}^{*}\widetilde{W}_{1}x_{0,j,i}+b_{1}^{*}+\widetilde{b}_{1}$ (5.58)
$=$ $\underbrace{W_{1}^{*}x_{0,j,i}+b_{1}^{*}}_{\in B_{\rho}(\beta_{1}u_{Q})}+W_{1}^{*}(\widetilde{W}_{1}-{\bf 1}_{Q\times Q})x_{0,j,i}+\widetilde{b}_{1}$
where
$|W_{1}^{*}(\widetilde{W}_{1}-{\bf 1}_{Q\times Q})x_{0,j,i}+\widetilde{b}_{1}|$$\leq$ $\|W_{1}^{*}\|_{op}\|\widetilde{W}_{1}-{\bf 1}_{Q\times Q}\|_{op}|x_{0,j,i}|+|\widetilde{b}_{1}|$ (5.59)
$\leq$ $\|W_{1}^{*}\|_{op}\;\epsilon\;\max_{j,i}|x_{0,j,i}|+\epsilon$
$\leq$ $C\epsilon$
for some constant $C$ independent of $j,i$. Therefore, we conclude that (5.58) is contained in $B_{\rho+C\epsilon}(\beta_{1}u_{Q})$, which lies in the positive sector ${\mathbb{R}}_{+}^{Q}\subset{\mathbb{R}}^{Q}$ for sufficiently small $\epsilon>0$. But this implies that
$\sigma(W_{1}^{*}\widetilde{W}_{1}X_{0}+B_{1}^{*}+\widetilde{B}_{1})=W_{1}^{*}\widetilde{W}_{1}X_{0}+B_{1}^{*}+\widetilde{B}_{1}\,.$ (5.60)
Thus, all arguments leading to the result in (3.29) stating that
${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]=\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}\,,$ (5.61)
for the choice of weights and biases $W_{i}^{*},b_{i}^{*}$, equally apply to the infinitesimally deformed $W_{1}^{*}\widetilde{W}_{1}$, $b_{1}^{*}+\widetilde{b}_{1}$ and the corresponding expressions for $W_{2}^{*}\widetilde{W}_{2}$, $b_{2}^{*}+\widetilde{b}_{2}$.
We now make the crucial observation that the right hand side of (5.61) does not depend on $W_{i}^{*},b_{i}^{*}$, as ${\mathcal{P}}^{\perp}$ only depends on the training inputs $X_{0}$; see (5.26).
Therefore,
${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]={\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*}\widetilde{W}_{i},b_{i}^{*}+\widetilde{b}_{i}]\,,$ (5.62)
and the variation in the perturbations $\widetilde{W}_{i},\widetilde{b}_{i}$ is zero. This implies that ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is a local minimum of the cost function, as it was obtained from a least square minimization problem. In particular, it is degenerate, and by repeating the above argument, one concludes that it has the same value for all weights and biases $W_{i},b_{i}$ which allow for the condition (5.55) to be satisfied.
Finally, we note that the reparametrization of the training inputs $X_{0}\rightarrow KX_{0}$, for any arbitrary $K\in GL(Q)$, induces
$\overline{X_{0}}\rightarrow K\overline{X_{0}}\;\;\;,\;\;\Delta X_{0}\rightarrow K\Delta X_{0}\,.$ (5.63)
In particular, this implies that
$\Delta_{1}^{rel}$ $=$ $(\overline{X_{0}^{red}})^{-1}\Delta X_{0}$ (5.64)
$\rightarrow$ $(K\overline{X_{0}^{red}})^{-1}K\Delta X_{0}$
$=$ $(\overline{X_{0}^{red}})^{-1}K^{-1}K\Delta X_{0}$
$=$ $\Delta_{1}^{rel}$
is invariant under this reparametrization, and hence,
$\Delta_{2}^{rel}$ $=$ $\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}$ (5.65)
also is. From the third and fourth lines in (5.44) follows that ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is a function only of $\Delta_{2}^{rel}$. We note that the reparametrization invariance of ${\mathcal{P}}$ (and thus of ${\mathcal{P}}^{\perp}$) can also be directly verified from (5.26) by inspection. This implies that ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is invariant under reparametrizations of the training inputs $X_{0}\rightarrow KX_{0}$, for all $K\in GL(Q)$.
This concludes the proof of Theorem 3.2. ∎
6. Constructive training of shallow network
In this section, we follow the proof of Theorem 3.1 in order to determine a constructively trained shallow network, which does not invoke the gradient descent algorithm.
In our approach, all relevant operations are confined to a $Q\leq M$ dimensional linear subspace of ${\mathbb{R}}^{M}$. To this end, we calculate
$\overline{X_{0}^{red}}:=[\overline{x_{0,1}}\cdots\overline{x_{0,Q}}]\in{\mathbb{R}}^{M\times Q}\,.$ (6.1)
We determine the Penrose inverse
${\rm Pen}[\overline{X_{0}^{red}}]=\big{(}(\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}}\big{)}^{-1}(\overline{X_{0}^{red}})^{T}\in{\mathbb{R}}^{Q\times M}$ (6.2)
and the orthoprojector
$P:=\overline{X_{0}^{red}}{\rm Pen}[\overline{X_{0}^{red}}]\in{\mathbb{R}}^{M\times M}\,.$ (6.3)
Subsequently, we determine $W_{1}=R\in O(M)$ which acts by rotating the ranges of $P$ and $P^{\perp}$ to be in alignment with the coordinate vectors $e_{j}:=(0,\dots 0,1,0,\dots,0)$. As $P$ is symmetric, this amounts to solving the diagonalization problem
$P=R^{T}P_{R}R\,.$ (6.4)
$P_{R}$ is diagonal, with $Q$ diagonal entries equaling $1$, and all other $M^{2}-Q$ matrix entries being 00. The orthogonal matrix $R$ diagonalizing $P$ also diagonalizes $P^{\perp}={\bf 1}-P$. We let
$W_{1}^{*}=R\in O(M)\,.$ (6.5)
Moreover, we set
$W_{2}^{*}=\widetilde{W}_{2}R^{T}$ (6.6)
where we calculate
$\widetilde{W}_{2}=Y((\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}})^{-1}(\overline{X_{0}^{red}})^{T}\in{\mathbb{R}}^{Q\times M}\,.$ (6.7)
To determine $B_{1}^{*}$, we split $b_{1}^{*}=P_{R}b_{1}^{*}+P^{\perp}_{R}b_{1}^{*}$, and set
$P_{R}b_{1}^{*}=\beta_{1}P_{R}u_{M}\,,$ (6.8)
recalling $u_{M}=(1,1,\dots,1)^{T}\in{\mathbb{R}}^{M}$ from (4.25). We choose $\beta_{1}\geq 0$ large enough to ensure that $RPX_{0}+P_{R}B_{1}^{*}$ is component-wise strictly positive. A sufficient condition is $\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|$. Moreover, we set
$P_{R}^{\perp}b_{1}^{*}=-\delta P^{\perp}_{R}u_{M}\,$ (6.9)
with $\delta$ as defined in (3.8). This defines $B_{1}^{*}=[b_{1}^{*}\cdots b_{1}^{*}]\in{\mathbb{R}}^{M\times N}$. Finally, we let
$b_{2}^{*}=-W_{2}^{*}P_{R}b_{1}^{*}\,.$ (6.10)
The set $W_{1}^{*},W_{2}^{*},b_{1}^{*},b_{2}^{*}$ defines the constructively trained shallow network. No use of the gradient descent algorithm is required.
7. Proof of Theorem 3.3
Here, we prove Theorem 3.3, and elucidate the geometric meaning of the constructively trained network obtained in Section 6. Given a test input $x\in{\mathbb{R}}^{M}$, the constructively trained network metrizes the linear subspace ${\rm range}(P)$ of the input space ${\mathbb{R}}^{M}$, and determines to which equivalence class of training inputs $Px$ is closest. To elucidate this interpretation, we observe that (6.7) implies that
$\widetilde{W}_{2}P=\widetilde{W}_{2}\;\;\;{\rm and}\;W_{2}^{*}P_{R}=W_{2}^{*}\,$ (7.1)
and hence,
$\widetilde{W}_{2}\overline{X_{0}^{red}}=Y\,,$ (7.2)
or in terms of column vectors,
$W_{2}^{*}RP\overline{x_{0,j}}=\widetilde{W}_{2}\overline{x_{0,j}}=y_{j}\,.$ (7.3)
Therefore,
${\mathcal{C}}_{j}[x]$ $=$ $|W_{2}^{*}(W_{1}^{*}x+b_{1}^{*})_{+}+b_{2}^{*}-y_{j}|$ (7.4)
$=$ $|W_{2}^{*}(Rx+b_{1}^{*})_{+}+b_{2}^{*}-y_{j}|$
$=$ $|W_{2}^{*}(RPx+RP^{\perp}x+P_{R}b_{1}^{*}+P^{\perp}_{R}b_{1}^{*})_{+}+b_{2}^{*}-y_{j}|$
$=$ $|W_{2}^{*}((P_{R}RPx+P_{R}b_{1}^{*})_{+}+(P^{\perp}_{R}RP^{\perp}x+P^{\perp}_{R}b_{1}^{*})_{+})+b_{2}^{*}-y_{j}|$
$=$ $|W_{2}^{*}(P_{R}RPx-y_{j})+W_{2}(P^{\perp}_{R}RP^{\perp}x-\delta P^{\perp}_{R}u)_{+}|$
$=$ $|W_{2}^{*}(P_{R}RP(x-\overline{x_{0,j}}))|$
where by construction of $P$, we have $P\overline{x_{0,j}}=\overline{x_{0,j}}$. Passing to the fourth line, we used $P_{R}RP=RP$, passing to the fifth line, we recalled $W_{2}^{*}P_{R}b_{1}^{*}+b_{2}^{*}=0$, and passing to the sixth line, we used (7.3). Moreover, we recall that $u$ is defined in (4.25). To pass to the sixth line, we also used that
$P^{\perp}_{R}(P^{\perp}_{R}RP^{\perp}x-\delta P^{\perp}_{R}u)_{+}=(P^{\perp}_{R}RP^{\perp}x-\delta P^{\perp}_{R}u)_{+}$ (7.5)
because the vector inside $(\cdot)_{+}$ only has components in the range of $P^{\perp}_{R}$, which is diagonal in the given coordinate system, and because $(\cdot)_{+}$ acts component-wise. Therefore, since also $W_{2}^{*}P_{R}=W_{2}^{*}$, we have
$W_{2}^{*}(P^{\perp}_{R}RP^{\perp}x-\delta P^{\perp}_{R}u)_{+}$ (7.6)
$=$ $W_{2}^{*}P_{R}P^{\perp}_{R}(P^{\perp}_{R}RP^{\perp}x-\delta P^{\perp}_{R}u)_{+}$
$=$ $0\,.$
With $W_{2}^{*}P_{R}RP=\widetilde{W}_{2}P$, (7.4) implies that
${\mathcal{C}}_{j}[x]=d_{\widetilde{W}_{2}}(Px,\overline{x_{0,j}})$ (7.7)
where
$d_{\widetilde{W}_{2}}(x,y):=|\widetilde{W}_{2}P(x-y)|\,,$ (7.8)
for $x,y\in{\rm range}(P)$, defines a metric in the range of $P$, which is a $Q$-dimensional linear subspace of the input space ${\mathbb{R}}^{M}$. This is because the matrix $\widetilde{W}_{2}=\widetilde{W}_{2}P$ has full rank $Q$.
We can therefore reformulate the identification of an input $x\in{\mathbb{R}}^{M}$ with an output $y_{j^{*}}$ via the constructively trained shallow network as the solution of the metric minimization problem
$j^{*}={\rm argmin}_{j\in\{1,\dots,Q\}}(d_{\widetilde{W}_{2}}(Px,\overline{x_{0,j}}))$ (7.9)
in the range of $P$. ∎
Acknowledgments: T.C. gratefully acknowledges support by the NSF through the grant DMS-2009800, and the RTG Grant DMS-1840314 - Analysis of PDE. P.M.E. was supported by NSF grant DMS-2009800 through T.C.