### v4

# arXiv:2309.10370v4[cs.LG] 26 Feb 2026

Abstract.
In this paper, we approach the problem of cost (loss) minimization in underparametrized shallow ReLU networks through the explicit construction of upper bounds which appeal to the structure of classification data, without use of gradient descent. A key focus is on elucidating the geometric structure of approximate and precise minimizers. We consider an ${\mathcal{L}}^{2}$ cost function, input space ${\mathbb{R}}^{M}$, output space ${\mathbb{R}}^{Q}$ with $Q\leq M$, and training input sample size that can be arbitrarily large. We prove an upper bound on the minimum of the cost function of order $O(\delta_{P})$ where $\delta_{P}$ measures the signal-to-noise ratio of training data. In the special case $M=Q$, we explicitly determine an exact degenerate local minimum of the cost function, and show that the sharp value differs from the upper bound obtained for $Q\leq M$ by a relative error $O(\delta_{P}^{2})$. The proof of the upper bound yields a constructively trained network; we show that it metrizes a particular $Q$-dimensional subspace in the input space ${\mathbb{R}}^{M}$. We comment on the characterization of the global minimum of the cost function in the given context.
1. Introduction
Applications of neural networks have rapidly become indispensable in a vast range of research disciplines, and their technological impact in recent years has been enormous. However, despite those successes, the fundamental conceptual reasons underlying their functioning are currently insufficiently understood, and remain the subject of intense investigation. The common approach to the minimization of the cost function is based on the gradient flow generated by it, and a variety of powerful algorithms have been developed for this purpose that can be highly successful in suitable application scenarios, see e.g. [LBH15, Aro18, GK23]; nevertheless, even if a global or sufficiently good local minimum is obtained, the properties of the minimizing parameters (weights, biases) in general remain quite mysterious, see e.g., [HR18, NRS21].
For linear neural networks and the $\mathcal{L}^{2}$ cost, critical points are well understood. Consider a matrix $X\in\mathbb{R}^{M\times N}$, whose columns are made up from the vectors in a training data set, and a matrix $Y\in\mathbb{R}^{Q\times N}$ whose columns are the corresponding target vectors. It is well known that a solution to
$\min_{W}\|WX-Y\|_{\mathcal{L}^{2}}$ (1.1)
is given by $W=Y\,{\rm Pen}[X]$, where ${\rm Pen}[X]$ is the Penrose inverse or pseudoinverse of $X$.
It is helpful to think in terms of projections. Note that ${\rm Pen}[X]X$ is a projector onto the orthogonal complement of the kernel of $X$, ${\rm Pen}[X]X=\mathcal{P}_{(\ker X)^{\perp}}$. Therefore, another way of stating the solution to the linear regression problem is
$\min_{W}\|WX-Y\|_{\mathcal{L}^{2}}=\left\lVert{Y{\rm Pen}[X]X-Y}\right\rVert_{\mathcal{L}^{2}}=\|Y\mathcal{P}_{(\ker X)^{\perp}}^{\perp}\|_{\mathcal{L}^{2}}.$ (1.2)
One can bound this as follows
$\left\lVert{Y\mathcal{P}_{\ker X}}\right\rVert_{\mathcal{L}^{2}}\leq\left\lVert{Y}\right\rVert_{op}\sqrt{\dim\ker X}=\left\lVert{Y}\right\rVert_{op}\sqrt{N-{\rm rank}X}.$ (1.3)
In the overdetermined case $M\geq N$, when $X$ is full-rank the cost can be taken to zero even with a linear network. We assume here that $N$ can be arbitrarily large, so that a zero-loss global minimum is not guaranteed by (1.3).
One of the goals of introducing non-linearities is to increase the complexity of the model. In this work, we study shallow neural networks
$X^{(2)}=W_{2}X^{(1)}+B_{2}\in\mathbb{R}^{Q\times N},\quad X^{(1)}=\sigma(W_{1}X+B_{1})\in\mathbb{R}^{M\times N},$ (1.4)
where $\sigma$ is the source of non-linearity, called an activation function. The ReLU activation function is a popular choice in applications; it is defined on vectors component-wise as the ramp function $(a)_{+}=\max\{0,a\}$. Because it acts as the identity function on vectors with non-negative components, it is convenient to study subcases where the network becomes linear so that results for linear networks apply, and this has been the subject of a few works [Kaw16, ZL18].
Minimizing the $\mathcal{L}^{2}$ cost with respect to $W_{2}$ yields
$\min_{W_{2}}\left\lVert{W_{2}X^{(1)}+B_{2}-Y}\right\rVert_{\mathcal{L}^{2}}$ $=\left\lVert{(Y-B_{2}){\rm Pen}[X^{(1)}]X^{(1)}-(Y-B_{2})}\right\rVert_{\mathcal{L}^{2}}$
$=\left\lVert{(Y-B_{2})\mathcal{P}_{\ker X^{(1)}}}\right\rVert_{\mathcal{L}^{2}}.$ (1.5)
The rank of $\mathcal{P}_{\ker X^{(1)}}$ is influenced by the coordinates picked out by $W_{1}$ and the ReLU activation function. When working with classification, there is a natural choice of coordinate frame: Suppose the set of training data is split into $Q$ classes. Then we may consider the mean of each class, and these vectors form the columns of a matrix $\overline{X_{0}^{red}}\in\mathbb{R}^{M\times Q}$. The deviation of each vector in the training data set from the corresponding target class mean is stored in the matrix $\Delta X_{0}$. In Theorem 3.1, we derive an upper bound on the minimum of the cost function with an explicit construction of suitable weights and biases:
$\min_{W_{1},b_{1},W_{2},b_{2}}\left\lVert{W_{2}X^{(1)}+B_{2}-Y}\right\rVert_{\mathcal{L}^{2}}$ $\leq C\,\|Y{\rm Pen}[\overline{X_{0}^{red}}]\Delta X_{0}\|_{\mathcal{L}^{2}}.$ (1.6)
Define the parameter
$\delta_{P}:=\sup_{\Delta x_{0}}\left\lvert{{\rm Pen}[\overline{X_{0}^{red}}]\Delta x_{0}}\right\rvert$, (1.7)
which measures the relative size between the averages in $\overline{X_{0}^{red}}$ and the deviations in $\Delta X_{0}$. Then (1.6) can be bounded in terms of this signal-to-noise ratio,
$\min_{W_{1},b_{1},W_{2},b_{2}}\left\lVert{W_{2}X^{(1)}+B_{2}-Y}\right\rVert_{\mathcal{L}^{2}}$ $\leq C\,\|Y\|_{op}\,\delta_{P}.$ (1.8)
The proof makes crucial use of the bias terms, where $b_{1}$ is constructed so that the terms in the span of the averages survive $\sigma$, and the terms in the complement disappear. When analyzing the particular case $M=Q$, the same strategy creates parameters which result in a linear neural network. Nevertheless, it is still possible to extract a meaningful bound: in Theorem 3.2, we explicitly determine an exact degenerate local minimum of the cost function, and show that the sharp value differs from the upper bound (1.6) obtained for $Q\leq M$ by a relative error $O(\delta_{P}^{2})$.
In Theorem 3.3, we use the parameters given in Theorem 3.1 to define a metric in (a subspace of) input space, and show that the problem of classifying multiclass data is equivalent to a metric minimization problem.
In Theorem 3.5, we prove a result for $Q=M$ closely related to Theorem 3.2, but circumventing the need for the network to be linear, via the introduction of a truncation map.
In Section 4, we compare our theoretical bounds with numerical results based on randomly initialized neural networks trained with gradient descent on synthetic data. We show that the bound improves with the increase of clustering of the data, whereby $\delta_{P}$ decreases; this matches our prediction.
Finally, we comment on the assumptions: We study shallow neural networks with dimensions of input, hidden and output layer $(M,M,Q)$, respectively, where $M\geq Q$ and at one point specialize to $M=Q$. While this is restrictive in the context of general neural networks, this can be useful as the study of the last layers of a deep network.
Moreover, we place no restriction on the number of training data points $N$ compared with the number of parameters, and so the results here hold for both underparametrized and overparametrized networks. In the strongly overparametrized regime $M>N$, every data point is its own cluster, so that the cluster variances are identically zero. Therefore, our construction yields zero training loss. This situation and its extension to deep networks is studied in detail in [BCC+25].
1.1. Related work
Early work on the loss landscape of neural networks already focused on shallow architectures. For linear networks, [BH89] analyzed the square loss for one-hidden-layer linear networks and showed that, under mild assumptions, every local minimum is global. [Kaw16] extended this to deep linear networks, proving the absence of bad local minima and characterizing saddle points, and subsequent works refined the description of critical points and their geometry, including explicit parametrizations and second-order analyses [ZL18, TKB20, AMG24]. For shallow ReLU networks, the loss landscape can be partly reduced to a linear problem when biases are ignored, and this reduction was exploited in [Kaw16, ZL18]. [Coo21] instead made crucial use of bias vectors to construct global minima that lie outside a fixed linear region for overparameterized networks, thereby showing that genuinely nonlinear structure in the loss landscape is already present in the shallow case.
Despite the dominance of very deep architectures in applications, shallow networks remain a central object in contemporary theory. From the optimization and generalization side, mean-field analyses for two-layer networks describe gradient-based training via limiting partial differential equations and yield convergence and generalization guarantees, e.g. [MMN18, MMM19]. In parallel, a substantial literature studies the implicit bias of gradient descent and gradient flow, both in regression and classification, e.g. [SHN+18, GLSS18, NCS19, LL20, JM23].
Shallow models also appear implicitly when one “freezes” all but the last few layers of a deep network. This perspective underlies work on linear classifier probes, where linear predictors are trained on top of intermediate representations to study how separability emerges across layers [AB17], and it is also closely connected to the unconstrained features model (UFM) and the phenomenon of neural collapse. In the UFM, one replaces the learned features by free variables and optimizes the last-layer problem jointly over logits and features, reducing a deep model to a structured shallow one [ZDZ+21, EW22, MPP22, TB22, HL24]. In all these settings, shallow architectures serve both as analytically tractable proxies and as faithful models of the final stages of deep networks.
Finally, shallow networks are also a minimal setting in which one can contrast lazy-training (kernel-like) and feature-learning regimes. In the infinite-width limit with appropriate scaling, gradient descent on two-layer ReLU networks evolves essentially as a linear model around initialization and leaves the representation almost fixed, the so-called lazy-training regime [JGH18, LXS+19, COB19]. By changing width, initialization scale, or training time one exits this regime and enters a genuinely nonlinear feature-learning regime where the representation itself evolves during training [LXMZ21]. Together with mean-field analyses that capture representation dynamics, this shows that shallow ReLU networks alone realize both extremes and the interpolation between them [MU25], while remaining simple enough that their loss landscape and implicit bias can be analyzed in detail.
1.2. Related work by the authors
We present the explicit construction of local and global ${\mathcal{L}}^{2}$ cost minimizers in deep neural networks for clustered data in a separate work [CE23], which invokes results developed in the paper at hand, combined with inspiration drawn from rigorous renormalization group analysis in quantum field theory, as for instance in [BCFS03]. In [CME25], we extend this construction to sequentially linearly separable data sets which are significantly less structured than in [CE23].
In Section 3.4, we address the structure of critical sets for shallow networks. More generally, the comparison of constructive training error bounds with those obtained from gradient descent is an interesting theoretical problem. A detailed analysis of the dynamical systems aspects of gradient descent and its geometric interpretation for both shallow and deep networks is given in [Che23, CME24].
Remark 1.1.
The inspiration for our approach is drawn from methods in mathematical physics aimed at determining the ground state energy for complex manybody quantum systems (such as large molecules), see for instance [LS10].
1.3. Outline of paper
This paper is organized as follows. In Section 2, we give a detailed description of the mathematical model and introduce notation. In Section 3, we present the main results, namely Theorem 3.1, which we prove in Section 5, Theorem 3.2, which we prove in Section 6, Theorem 3.3, which we prove in Section 7, and Theorem 3.5. In Section 4, we present some numerical experiments.
2. Definition of the mathematical model
In this section, we give a detailed introduction of the mathematical model in discussion, consisting of a shallow neural network for a classification task with $Q$ classes. We also define relevant notation.
We introduce weight matrices
$W_{1}\in{\mathbb{R}}^{M\times M}\;\;,\;\;W_{2}\in{\mathbb{R}}^{Q\times M}$ (2.1)
and bias matrices
$B_{1}$ $:=$ $b_{1}u_{N}^{T}=[b_{1}\cdots b_{1}]\in{\mathbb{R}}^{M\times N}$
$B_{2}$ $:=$ $b_{2}u_{N}^{T}=[b_{2}\cdots b_{2}]\in{\mathbb{R}}^{Q\times N}$ (2.2)
where $b_{1}\in{\mathbb{R}}^{M}$ and $b_{2}\in{\mathbb{R}}^{Q}$ are column vectors, and
$u_{N}=(1,1,\dots,1)^{T}\;\;\in{\mathbb{R}}^{N}\,.$ (2.3)
Moreover, we choose to consider the ReLU activation function
$\sigma:{\mathbb{R}}^{M\times N}$ $\rightarrow$ ${\mathbb{R}}_{+}^{M\times N}$
$A=[a_{ij}]$ $\mapsto$ $[(a_{ij})_{+}]$ (2.4)
where
$(a)_{+}:=\max\{0,a\}$ (2.5)
is the standard ramp function. Then we define the neural network as follows: for a matrix of inputs $X_{0}\in\mathbb{R}^{M\times N}$,
$X^{(1)}:=\sigma(W_{1}X_{0}+B_{1})\in{\mathbb{R}}_{+}^{M\times N}$ (2.6)
is the hidden layer, and
$X^{(2)}:=W_{2}X^{(1)}+B_{2}\in{\mathbb{R}}^{Q\times N}\,$ (2.7)
is the terminal layer.
As the labels for the $Q$ classes, we consider a family $\{y_{j}\}_{j=1}^{Q}$ of linearly independent target vectors, and define a target matrix
$Y:=[y_{1},\dots,y_{Q}]\in{\mathbb{R}}^{Q\times Q},$ (2.8)
which is invertible. We also define the extension of the matrix of labels, $Y^{ext}$, by
$Y^{ext}:=[Y_{1}\cdots Y_{Q}]\in{\mathbb{R}}^{Q\times N}$ (2.9)
where
$Y_{j}:=[y_{j}\cdots y_{j}]\in{\mathbb{R}}^{Q\times N_{j}}$ (2.10)
with $N_{j}$ copies of the same target column vector $y_{j}$. Clearly, $Y^{ext}$ has full rank $Q$.
We let ${\mathbb{R}}^{M}$ denote the input space with coordinate system defined by the orthonormal basis vectors $\{e_{\ell}=(0,\dots,0,1,0,\dots,0)^{T}\in{\mathbb{R}}^{M}\,|\,\ell=1,\dots,M\}$. For $j\in\{1,\dots,Q\}$, and $N_{j}\in{\mathbb{N}}$,
$x_{0,j,i}\in{\mathbb{R}}^{M}\;\;,\;\;i\in\{1,\dots,N_{j}\}$ (2.11)
denotes the $i$-th training input vector corresponding to the $j$-th target vector $y_{j}$. We define the matrix of all training inputs belonging to $y_{j}$ by
$X_{0,j}:=[x_{0,j,1}\cdots x_{0,j,i}\cdots x_{0,j,N_{j}}]\,.$ (2.12)
Then, we define the matrix of training inputs
$X_{0}:=[X_{0,1}\cdots X_{0,j}\cdots X_{0,Q}]\in{\mathbb{R}}^{M\times N}$ (2.13)
where $N:=\sum_{j=1}^{Q}N_{j}$.
Let
$\overline{x_{0,j}}:=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}x_{0,j,i}\in{\mathbb{R}}^{M}$ (2.14)
denote the average over the equivalence class of all training input vectors corresponding to the $j$-th class, and
$\Delta x_{0,j,i}:=x_{0,j,i}-\overline{x_{0,j}}\,.$ (2.15)
Moreover, let
$\overline{X_{0,j}}:=[\overline{x_{0,j}}\cdots\overline{x_{0,j}}]\in{\mathbb{R}}^{M\times N_{j}}$ (2.16)
and
$\overline{X_{0}}:=[\overline{X_{0,1}}\cdots\overline{X_{0,Q}}]\in{\mathbb{R}}^{M\times N}\,.$ (2.17)
Similarly, let
$\Delta X_{0,j}:=[\Delta x_{0,j,1}\cdots\Delta x_{0,j,N_{j}}]\in{\mathbb{R}}^{M\times N_{j}}\,.$ (2.18)
and
$\Delta X_{0}:=[\Delta X_{0,1}\cdots\Delta X_{0,Q}]\in{\mathbb{R}}^{M\times N}\,.$ (2.19)
Then,
$X_{0}=\overline{X_{0}}+\Delta X_{0}\,$, (2.20)
and we let
$\delta:=\sup_{j,i}|\Delta x_{0,j,i}|\ \,.$ (2.21)
Next, we define the reduction of $\overline{X_{0}}$ as
$\overline{X_{0}^{red}}:=[\overline{x_{0,1}}\cdots\overline{x_{0,Q}}]\in{\mathbb{R}}^{M\times Q}$ (2.22)
and assume that its rank is maximal, ${\rm rank}(\overline{X_{0}^{red}})=Q$. The latter condition is necessary for the averages of training inputs to be able to distinguish between the different target values $y_{j}$, $j=1,\dots,Q$. Moreover, in this case the matrix
$(\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}}\in{\mathbb{R}}^{Q\times Q}$ (2.23)
is invertible, and we can define the Penrose inverse [^1] Also called Moore–Penrose inverse or pseudoinverse and denoted $A^{+}$ instead of ${\rm Pen}[A]$. of $\overline{X_{0}^{red}}$,
${\rm Pen}[\overline{X_{0}^{red}}]:=\big((\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}}\big)^{-1}(\overline{X_{0}^{red}})^{T}\in{\mathbb{R}}^{Q\times M}.$ (2.24)
The orthoprojector onto the range of $\overline{X_{0}^{red}}$ is then given by
$P:=\overline{X_{0}^{red}}{\rm Pen}[\overline{X_{0}^{red}}]\in{\mathbb{R}}^{M\times M}$ (2.25)
and we note that ${\rm Pen}[\overline{X_{0}^{red}}]\overline{X_{0}^{red}}={\bf 1}_{Q\times Q}$. The projector property $P^{2}=P$ is thus easily verified, and orthogonality with respect to the Euclidean inner product $(\cdot,\cdot)$ on ${\mathbb{R}}^{M}$ holds due to $P^{T}=P$, whereby $(v,Pw)=(Pv,w)$ for any $v,w\in{\mathbb{R}}^{M}$. In particular, we have
$P\overline{X_{0}^{red}}=\overline{X_{0}^{red}}\;\;\;{\rm and}\;\;\;P\overline{X_{0}}=\overline{X_{0}}\,$, (2.26)
by construction.
As the objective function, we consider the ${\mathcal{L}}^{2}$ Schatten class (or Hilbert-Schmidt) cost function
${\mathcal{C}}[W_{j},b_{j}]$ $:=\frac{1}{\sqrt{N}}\|X^{(2)}-Y^{ext}\|_{{\mathcal{L}}^{2}}$ (2.27)
with
$\|A\|_{{\mathcal{L}}^{2}}\equiv\sqrt{{\rm Tr}(AA^{T})}\,$, (2.28)
where $A^{T}$ is the transpose of the matrix $A$. Equivalently,
$\mathcal{C}[W_{j},b_{j}]$ $=\sqrt{\frac{1}{N}\sum_{j=1}^{Q}\sum_{i=1}^{N_{j}}|W_{2}\sigma(W_{1}x_{0,j,i}+b_{1})+b_{2}-y_{j}|^{2}_{{\mathbb{R}}^{Q}}}.$ (2.29)
The cost is a function of the weights and biases, while the training inputs in $X_{0}$ are considered to be a fixed set of parameters. Training this shallow network amounts to finding the optimizers $W_{j}^{**},b_{j}^{**}$ for the minimization problem
${\mathcal{C}}[W_{j}^{**},b_{j}^{**}]=\min_{W_{j},b_{j}}{\mathcal{C}}[W_{j},b_{j}]\,.$ (2.30)
3. Statement of Main Results
In this section, we present the main results of this paper, an explicit upper bound on the cost function in Theorem 3.1 which leads to the construction of an optimized constructively trained shallow network, and its natural geometric interpretation in Theorem 3.3. In Theorem 3.2, we construct an explicit degenerate local minimum of the cost function, for the special case of input and output space having equal dimension, $M=Q$. In Theorem 3.5, we address the effects of truncation of the training input matrix due to the activation function, in the case $Q=M$.
All bounds derived here make use of the following quantity,
$\delta_{P}:=\sup_{i,j}\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,j,i}\right|\,$, (3.1)
which measures the relative size between $\overline{X_{0}^{red}}$ and $P\Delta X_{0}$, as $\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,j,i}\right|$ scales like the noise-to-signal ratio $\frac{|\Delta x|}{|x|}$ of training inputs. Note that it is scaling invariant under $X_{0}\rightarrow\lambda X_{0}$.
3.1. Upper bound on minimum of cost function for $M>Q$
In Theorem 3.1, we prove an upper bound on the minimum of the cost function with the following construction.
Let $R\in O(M)$ denote an orthogonal matrix that diagonalizes $P$, defined in (2.25). That is, the orthoprojectors
$P_{R}:=RPR^{T}\;\;\;{\rm and}\;P^{\perp}_{R}:=RP^{\perp}R^{T}$ (3.2)
are diagonal in the given coordinate system; this is important for compatibility with the fact that $\sigma$ acts component-wise. Namely, $R$ rotates the input data in such a way that ${\rm range}(P)$ is made to align with the coordinate axes; this decouples the action of $\sigma$ on the rotated ${\rm range}(P)$ from its action on the rotated ${\rm range}(P^{\perp})$. That is, $\sigma(v)=\sigma(P_{R}v)+\sigma(P^{\perp}_{R}v)$ for all $v\in{\mathbb{R}}^{M}$.
Letting
$u_{M}:=(1,1,\cdots,1)^{T}\in{\mathbb{R}}^{M}$ (3.3)
we may choose $\beta_{1}\geq 0$ sufficiently large ($\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|$ is sufficient), so that the projected, translated, and rotated training input vectors $RPx_{0,j,i}+\beta_{1}u_{M}$ are component-wise non-negative for all $i=1,\dots,N_{j}$, and $j=1,\dots,Q$.
Then, we construct an upper bound on the cost function by use of the following weights and biases. We choose
$W_{1}^{*}=R$ (3.4)
and $b_{1}^{*}=P_{R}b_{1}^{*}+P_{R}^{\perp}b_{1}^{*}$ where
$P_{R}b_{1}^{*}=\beta_{1}P_{R}u_{Q}\;\;\;,\;\;\;P^{\perp}_{R}b_{1}^{*}=-\delta P^{\perp}_{R}u_{M}\,$, (3.5)
with $\beta_{1}\geq 0$ large enough to ensure component-wise non-negativity of $RPX_{0}+P_{R}B_{1}^{*}$, where $B_{1}^{*}=b_{1}^{*}u_{N}^{T}$, so that
$\sigma(RPX_{0}+P_{R}B_{1}^{*})=RPX_{0}+P_{R}B_{1}^{*}\,.$ (3.6)
On the other hand, the MQM-Q rows of $RP^{\perp}X_{0}+P_{R}^{\perp}B_{1}^{*}$ are eliminated by way of
$\sigma(RP^{\perp}X_{0}+P_{R}^{\perp}B_{1}^{*})=0\,.$ (3.7)
We thus obtain a reduction of the dimension of the input space from $M$ to $Q$.
Passing to the output layer, we require $W_{2}^{*}$ to solve
$W_{2}^{*}R\overline{X_{0}^{red}}=Y\,$, (3.8)
which yields
$W_{2}^{*}=Y\;{\rm Pen}[\overline{X_{0}^{red}}]PR^{T}\,.$ (3.9)
Finally, we find that
$b_{2}^{*}=-W_{2}^{*}P_{R}b_{1}^{*}\,$, (3.10)
that is, it reverts the translation by $P_{R}b_{1}^{*}$ in the previous layer.
Theorem 3.1.
Let $Q\leq M\leq QM$. Assume that $R\in O(M)$ diagonalizes $P,P^{\perp}$, and let $\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|$.
Let ${\mathcal{C}}[W_{i}^{*},b_{i}^{*}]$ be the cost function evaluated for the trained shallow network defined by the following weights and biases,
$W_{1}^{*}=R\,$, (3.11)
and
$W_{2}^{*}=\widetilde{W}_{2}R^{T}\;\;\;{\rm where}\;\widetilde{W}_{2}=Y\;{\rm Pen}[\overline{X_{0}^{red}}]P\,.$ (3.12)
Moreover, $b_{1}^{*}=P_{R}b_{1}^{*}+P_{R}^{\perp}b_{1}^{*}$ with
$P_{R}b_{1}^{*}=\beta_{1}u_{M}\;\;,\;P_{R}^{\perp}b_{1}^{*}=-\delta P_{R}^{\perp}u_{M}\;\;\in{\mathbb{R}}^{M}\,$, (3.13)
for $u_{M}\in{\mathbb{R}}^{M}$ as in (3.3), and
$b_{2}^{*}=-W_{2}^{*}P_{R}b_{1}^{*}\;\;\in{\mathbb{R}}^{Q}\,.$ (3.14)
Then, the minimum of the cost function satisfies the upper bound
$\min_{W_{j},b_{j}}{\mathcal{C}}[W_{j},b_{j}]\leq{\mathcal{C}}[W_{i}^{*},b_{i}^{*}]\leq\frac{1}{\sqrt{N}}\|Y\;{\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}\|_{{\mathcal{L}}^{2}}\,$, (3.15)
which implies
$\min_{W_{j},b_{j}}{\mathcal{C}}[W_{j},b_{j}]\;\leq\;\|Y\|_{op}\;\delta_{P}\;$ (3.16)
where $N=\sum_{\ell=1}^{Q}N_{\ell}$.
A geometrically intuitive way of thinking about this construction is that $W_{1}^{*}=R$ orients the training input data with respect to the given coordinate system, in order to align it with the component-wise action of $\sigma$. This allows for a maximal rank reduction via $\sigma$, whereby the maximal possible amount of insignificant information is eliminated: $P_{R}b_{1}^{*}$ pulls the significant information (in the range of $P_{R}$) out of the kernel of $\sigma$, while $P_{R}^{\perp}b_{1}^{*}$ pushes the insignificant information (in the range of $P_{R}^{\perp}$) into the kernel of $\sigma$ whereby the latter is eliminated. Subsequently, $b_{2}^{*}$ places the significant information back into its original position, and $W_{2}^{*}$ matches it to the output matrix $Y$ in the sense of least squares, with respect to the ${\mathcal{L}}^{2}$-norm.
We observe that, for any $\lambda>0$, the rescaling of training inputs
$x_{0,j,i}\rightarrow\lambda x_{0,j,i}\;\;\forall j=1,\dots,Q\;\;,\;i=1,\dots,N_{j}\,$, (3.17)
induces
${\rm Pen}[\overline{X_{0}^{red}}]\rightarrow\lambda^{-1}{\rm Pen}[\overline{X_{0}^{red}}]\;\;\;,\;\Delta X_{0}\rightarrow\lambda\Delta X_{0}\,.$ (3.18)
Therefore, the upper bound in (3.15) is scaling invariant, as it only depends on the signal-to-noise ratio of training input data, which is controlled by $\delta_{P}$.
3.2. Exact degenerate local minimum in the case $M=Q$
We explicitly determine an exact local degenerate minimum of the cost function in the case $Q=M$ where the input and output spaces have the same dimension; here, we show that the upper bound obtained above differs from the sharp value by a relative error of order $O(\delta_{P}^{2})$. We note that in this situation, the projector defined in (2.25) satisfies $P={\bf 1}_{Q\times Q}$ and $P^{\perp}=0$. Moreover, $\overline{X_{0}^{red}}$ is invertible, and therefore, ${\rm Pen}[\overline{X_{0}^{red}}]=(\overline{X_{0}^{red}})^{-1}$.
We prove this result relative to a weighted variant of the cost function defined as follows. Let ${\mathcal{N}}\in{\mathbb{R}}^{N\times N}$ be the block diagonal matrix given by
${\mathcal{N}}:={\rm diag}(N_{j}{\bf 1}_{N_{j}\times N_{j}}\,|\,j=1,\dots,Q)\,.$ (3.19)
We introduce the inner product on ${\mathbb{R}}^{Q\times N}$
$(A,B)_{{\mathcal{L}}^{2}_{\mathcal{N}}}:={\rm Tr}(A{\mathcal{N}}^{-1}B^{T})$ (3.20)
and
$\|A\|_{{\mathcal{L}}^{2}_{\mathcal{N}}}:=\sqrt{(A,A)_{{\mathcal{L}}^{2}_{\mathcal{N}}}}\,.$ (3.21)
We define the weighted cost function
${\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]:=\|X^{(2)}-Y^{ext}\|_{{\mathcal{L}}^{2}_{\mathcal{N}}}\,.$ (3.22)
This is equivalent to
${\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]:=\sqrt{\sum_{j=1}^{Q}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}|W_{2}\sigma(W_{1}x_{0,j,i}+b_{1})+b_{2}-y_{j}|^{2}_{{\mathbb{R}}^{Q}}}\,.$ (3.23)
The weights $\frac{1}{N_{j}}$ ensure that the contributions to the cost function belonging to inputs $y_{j}$ do not depend on their sample sizes $N_{j}$. We note that for uniform sample sizes, where $\frac{N_{j}}{N}=\frac{1}{Q}$ $\forall j$, we have that ${\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]=\sqrt{Q}{\mathcal{C}}[W_{i},b_{i}]$.
Theorem 3.2.
Assume $M=Q<MQ$. Then, if ${\rm rank}(\overline{X_{0}})=Q$, let
${\mathcal{P}}:={\mathcal{N}}^{-1}X_{0}^{T}(X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}\;\;\in{\mathbb{R}}^{N\times N}\,.$ (3.24)
Its transpose ${\mathcal{P}}^{T}$ is an orthonormal projector in ${\mathbb{R}}^{N}$ with respect to the inner product $\langle u,v\rangle_{{\mathcal{N}}}:=(u,{\mathcal{N}}^{-1}v)$. In particular, $(A{\mathcal{P}},B)_{{\mathcal{L}}^{2}_{\mathcal{N}}}=(A,B{\mathcal{P}})_{{\mathcal{L}}^{2}_{\mathcal{N}}}$ for all $A,B\in{\mathbb{R}}^{Q\times N}$. Letting ${\mathcal{P}}^{\perp}={\bf 1}_{N\times N}-{\mathcal{P}}$, the weighted cost function satisfies the upper bound
$\min_{W_{j},b_{j}}{\mathcal{C}}_{{\mathcal{N}}}[W_{j},b_{j}]$ $\leq$ ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ (3.25)
$=$ $\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}$
$=$ $\big\|Y|\Delta_{2}^{rel}|^{\frac{1}{2}}\big(1+\Delta_{2}^{rel}\big)^{-\frac{1}{2}}\big\|_{{\mathcal{L}}^{2}}$
$\leq$ $(1-C_{0}\delta_{P}^{2})\;\|Y\;(\overline{X_{0}^{red}})^{-1}\Delta X_{0}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}\,$,
for a constant $C_{0}\geq 0$, and where
$\Delta_{2}^{rel}:=\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}\;\;,\;{\rm with}\;\Delta_{1}^{rel}:=(\overline{X_{0}^{red}})^{-1}\Delta X_{0}\,.$ (3.26)
The weights and biases realizing the upper bound are given by
$W_{1}^{*}={\bf 1}_{Q\times Q}$ (3.27)
and
$W_{2}^{*}=Y(\overline{X_{0}^{red}})^{T}(X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}\in{\mathbb{R}}^{Q\times Q}\,.$ (3.28)
Moreover,
$b_{1}^{*}=\beta_{1}u_{Q}\;\;,\;b_{2}^{*}=-W_{2}^{*}b_{1}^{*}\,$, (3.29)
with $u_{Q}\in{\mathbb{R}}^{Q}$ as in (3.3), and where $\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|$ so that
$\sigma(W_{1}^{*}X_{0}+B_{1}^{*})=W_{1}^{*}X_{0}+B_{1}^{*}\,.$ (3.30)
Notably,
$W_{2}^{*}=\widetilde{W}_{2}+O(\delta_{P}^{2})$ (3.31)
differs by $O(\delta_{P}^{2})$ from $\widetilde{W}_{2}$ in (3.12) used for the upper bound (3.15).
In particular, ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is a local minimum of the cost function; it is degenerate, assuming the same value for all $W_{i},b_{i}$ such that $W_{1},b_{1}$ satisfy the condition (3.30).
Moreover, ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is invariant under reparametrizations of the training inputs $X_{0}\rightarrow KX_{0}$, for all $K\in GL(Q)$.
The proof is given in Section 6.
3.3. Geometric interpretation
To match an arbitrary test input $x\in{\mathbb{R}}^{M}$ with one of the output vectors $y_{j}$, $j\in\{1,\dots,Q\}$, let for $x\in{\mathbb{R}}^{M}$,
${\mathcal{C}}_{j}[x]:=|W_{2}^{*}(W_{1}^{*}x+b_{1}^{*})_{+}+b_{2}^{*}-y_{j}|\,.$ (3.32)
Here, $(\cdot)_{+}$ denotes the ramp function (2.5), acting component-wise on a vector.
Given an arbitrary test input $x\in{\mathbb{R}}^{M}$,
$j^{*}={\rm argmin}_{j}{\mathcal{C}}_{j}[x]$ (3.33)
implies that $x$ matches the $j^{*}$-th output $y_{j^{*}}$.
The proof of Theorem 3.1 provides us with a particular set of weights $W_{1}^{*},W_{2}^{*}$ and biases $b_{1}^{*},b_{2}^{*}$ through an explicit construction, which yield the upper bound (3.15). We will refer to the shallow network trained with this choice of $W_{i}^{*},b_{i}^{*}$ as the constructively trained shallow network.
Theorem 3.3.
Assume $Q\leq M\leq QM$. Let $W_{i}^{*},b_{i}^{*}$, $i=1,2$, denote the weights and biases determined in Theorem 3.1. Let
$\widetilde{W}_{2}:=Y\;{\rm Pen}[\overline{X_{0}^{red}}]\in{\mathbb{R}}^{Q\times M}\,$, (3.34)
and define the metric
$d_{\widetilde{W}_{2}}(x,y):=|\widetilde{W}_{2}P(x-y)|\;\;\;{\rm for}\;x,y\in{\rm range}(P)$ (3.35)
on the $Q$-dimensional linear subspace ${\rm range}(P)\subset{\mathbb{R}}^{M}$, where $|\cdot|$ denotes the Euclidean norm on ${\mathbb{R}}^{Q}$. Then,
${\mathcal{C}}_{j}[x]=d_{\widetilde{W}_{2}}(Px,\overline{x_{0,j}})\,$, (3.36)
and matching an input $x\in{\mathbb{R}}^{M}$ with a target output $y_{j^{*}}$ via the constructively trained shallow network is equivalent to the solution of the metric minimization problem
$j^{*}={\rm argmin}_{j\in\{1,\dots,Q\}}(d_{\widetilde{W}_{2}}(Px,\overline{x_{0,j}}))$ (3.37)
on the range of $P$.
The proof is given in Section 7. The geometric interpretation of the construction is as follows.
First of all, we note that $\widetilde{W}_{2}P=\widetilde{W}_{2}$ has full rank $Q$, therefore $x\mapsto|\widetilde{W}_{2}Px|^{2}$ is a non-degenerate quadratic form on the range of $P$. Therefore, (3.35) indeed defines a metric on the range of $P$.
The constructively trained shallow network thus obtained matches a test input $x\in{\mathbb{R}}^{M}$ with an output vector $y_{j^{*}}$ by splitting
$x=Px+P^{\perp}x\,$, (3.38)
and by determining which of the average training input vectors $\overline{x_{0,j}}$, $j=1,\dots,Q$, is closest to $Px$ in the $d_{\widetilde{W}_{2}}$ metric. The trained shallow network cuts off the components $P^{\perp}x$.
3.4. Dependence on truncation
Our next result addresses the effect of truncations enacted by the activation function $\sigma$ in the case $M=Q$. For the results in Theorem 3.2, we assumed that $W_{1},b_{1}$ are in the region (depending on $X_{0}$) such that
$\sigma(W_{1}X_{0}+B_{1})=W_{1}X_{0}+B_{1}$ (3.39)
holds. Here we discuss the situation in which $\sigma$ acts nontrivially. In this case, we assume that $W_{1},b_{1}$ are given, with $W_{1}\in{\mathbb{R}}^{Q\times Q}$ invertible, and observe that all matrix components of
$X^{(1)}=\sigma(W_{1}X_{0}+B_{1})$ (3.40)
are non-negative, as $X^{(1)}$ lies in the image of $\sigma$, and hence,
$\sigma(X^{(1)})=X^{(1)}\,.$ (3.41)
We define the truncation map $\tau_{W_{1},b_{1}}$ as follows. It will play a major role in our explicit construction of local and global cost minimizers in deep neural networks in [CE23].
Definition 3.4.
Let $W_{1}\in GL(M)$, $b_{1}\in{\mathbb{R}}^{M}$. Then, the truncation map $\tau_{W_{1},b_{1}}:{\mathbb{R}}^{M\times N}\rightarrow{\mathbb{R}}^{M\times N}$ is defined by
$\tau_{W_{1},b_{1}}(X_{0})$ $:=$ $W_{1}^{-1}(\sigma(W_{1}X_{0}+B_{1})-B_{1})$ (3.42)
$=$ $W_{1}^{-1}(X^{(1)}-B_{1})\,.$
That is, $\tau_{W_{1},b_{1}}=a_{W_{1},b_{1}}^{-1}\circ\sigma\circ a_{W_{1},b_{1}}$ under the affine map $a_{W_{1},b_{1}}:X_{0}\mapsto W_{1}X_{0}+B_{1}$.
We say that $\tau_{W_{1},b_{1}}$ is rank preserving (with respect to $X_{0}$) if both
${\rm rank}(\tau_{W_{1},b_{1}}(X_{0}))$ $=$ ${\rm rank}(X_{0})$
${\rm rank}(\overline{\tau_{W_{1},b_{1}}(X_{0})})$ $=$ ${\rm rank}(\overline{X_{0}})$ (3.43)
hold, and that it is rank reducing otherwise.
Then, we verify that
$\sigma(W_{1}\tau_{W_{1},b_{1}}(X_{0})+B_{1})$ $=$ $\sigma(X^{(1)})=X^{(1)}$ (3.44)
$=$ $W_{1}\tau_{W_{1},b_{1}}(X_{0})+B_{1}\,.$
This means that while the condition (3.39) does not hold for $X_{0}$, it does hold for $\tau_{W_{1},b_{1}}(X_{0})$.
Accordingly, we can define the matrices
$\overline{\tau_{W_{1},b_{1}}(X_{0})}\;\;,\;\overline{(\tau_{W_{1},b_{1}}(X_{0}))^{red}}\;\;,\;\Delta(\tau_{W_{1},b_{1}}(X_{0}))$ (3.45)
in analogy to $\overline{X_{0}},\overline{X_{0}^{red}},\Delta X_{0}$.
If no rank reduction is induced by the truncation, $\overline{(\tau_{W_{1},b_{1}}(X_{0}))^{red}}$ is invertible, and we obtain the following theorem.
Theorem 3.5.
Let $Q=M$, and assume that the truncation map $\tau_{W_{1},b_{1}}(X_{0})$ is rank preserving. Then, for any fixed $(W_{1},b_{1})\in GL(Q)\times{\mathbb{R}}^{Q}$,
$\min_{W_{2},b_{2}}{\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]=\|Y^{ext}{\mathcal{P}}^{\perp}_{\tau_{W_{1},b_{1}}(X_{0})}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}$ (3.46)
where the projector ${\mathcal{P}}_{\tau_{W_{1},b_{1}}(X_{0})}$ is obtained from (3.24) by substituting the truncated training inputs $\tau_{W_{1},b_{1}}(X_{0})$ for $X_{0}$. In particular, one explicitly has
$\min_{W_{2},b_{2}}{\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]$ $=$ $\big\|Y|\Delta_{2}^{rel,tr}|^{\frac{1}{2}}\big(1+\Delta_{2}^{rel,tr}\big)^{-\frac{1}{2}}\big\|_{{\mathcal{L}}^{2}}$ (3.47)
$\leq$ $(1-C_{0}\delta_{P,tr}^{2})\|Y\Delta_{1}^{rel,tr}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}$
where
$\Delta_{2}^{rel,tr}$ $:=$ $\Delta_{1}^{rel,tr}{\mathcal{N}}^{-1}(\Delta_{1}^{rel,tr})^{T}\;\;,\;{\rm with}\;$
$\Delta_{1}^{rel,tr}$ $:=$ $(\overline{(\tau_{W_{1},b_{1}}(X_{0}))^{red}})^{-1}\Delta(\tau_{W_{1},b_{1}}(x_{0,j,i}))\,$, (3.48)
in analogy to (3.25), and for some $C_{0}\geq 0$, where
$\delta_{P,tr}:=\sup_{j,i}\left|(\overline{(\tau_{W_{1},b_{1}}(X_{0}))^{red}})^{-1}\Delta(\tau_{W_{1},b_{1}}(x_{0,j,i}))\right|$ (3.49)
measures the signal-to-noise ratio of the truncated training input data.
Proof.
Due to (3.44), we may follow the proof of Theorem 3.2 verbatim, and we find that for the given fixed choice of $W_{1},b_{1}$, minimization in $W_{2},b_{2}$ yields (3.46) and (3.47). ∎
In general, (3.47) does not constitute a stationary solution with respect to $W_{1},b_{1}$, as the right hand sides of (3.46), (3.47) generically depend nontrivially on $W_{1},b_{1}$. It thus remains to determine the infimum of (3.46), (3.47) with respect to $W_{1},b_{1}$, in order to find a candidate for the global minimum. The latter, however, might depend sensitively on the detailed properties of $\Delta X_{0}$, which are random.
On the other hand, in the situation addressed in Theorem 3.2, $W_{1},b_{1}$ are contained in the parameter region ${\mathcal{F}}_{X_{0}}$ (depending on $X_{0}$) defined by the fixed point relation for the truncation map (equivalent to (3.39)),
${\mathcal{F}}_{X_{0}}:=\{(W_{1},b_{1})\in GL(Q)\times{\mathbb{R}}^{Q}\;|\;\tau_{W_{1},b_{1}}(X_{0})=X_{0}\}\,.$ (3.50)
As a result, we indeed obtained a degenerate stationary solution with respect to $W_{1},b_{1}$,
$\min_{(W_{1},b_{1})\in{\mathcal{F}}_{X_{0}};W_{2},b_{2}}{\mathcal{C}}_{{\mathcal{N}}}[W_{i},b_{i}]=\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}$ (3.51)
in (3.25), with $\mathcal{P}$} given in (3.24). Hence, ${\mathcal{F}}_{X_{0}}$ parametrizes an invariant manifold of equilibria of the gradient descent flow, which corresponds to a level surface of the cost function.
Remark 3.6.
The global minimum of the cost function has either the form (3.46) for an optimizer $(W_{1}^{**},b_{1}^{**})$ (with a rank preserving truncation), or (3.51). This is because any other scenario employs a rank reducing truncation. Since $Y^{ext}$ has rank $Q$, minimization of (3.22) under the constraint ${\rm rank}(\tau_{W_{1},b_{1}}(X_{0}))<Q$ cannot yield a global minimum.
The key point in minimizing (3.46) with respect to $(W_{1},b_{1})$ is to determine a rank preserving truncation which minimizes the signal to noise ratio of the truncated training inputs. Whether for a local or global minimum of the form (3.46) or (3.51), respectively, a shallow network trained with the corresponding weights and biases $W_{i},b_{i}$, $i=1,2$, will be able to match inputs $x\in{\mathbb{R}}^{Q}$ to outputs $y_{j}$, through a straightforward generalization of Theorem 3.3.
We expect these considerations to carry over to the general case $Q<M$, but leave a detailed discussion to future work.
4. Experiments
To assess the validity of our theoretical results in practical training scenarios, we generated synthetic data using a Gaussian mixture model in $\mathbb{R}^{M}$ with $Q$ classes, each having randomly distributed means and varying standard deviations. For each dataset, constructed with a fixed standard deviation and $Q$ classes, we initialized and trained a ReLU neural network with architecture $(M,M,Q)$, as defined in (2.6)-(2.7), to classify the data according to the cost function (2.27). Each configuration was trained 100 times. Figures 1 and 2 display the average final cost across runs, the corresponding average initial cost, and the bound (3.15) from Theorem 3.1, which we repeat here:
$\text{Bound}=\frac{1}{\sqrt{N}}\left\lVert{\left(\overline{X_{0}^{red}}\right)^{+}\Delta X_{0}}\right\rVert_{\mathcal{L}^{2}}.$ (4.1)
Note that we taking the label vectors to be the canonical basis vectors in $\mathbb{R}^{Q}$, so that $Y$ is the identity.
Figure 1. (Average of) initial ($C_{init}$) and final ($C_{final}$) cost of randomly initialized neural network(s) trained to classify Gaussian mixture data with $Q$ classes, for different (fixed in each plot) cluster standard deviations, plotted against the bound (4.1) computed for each data set.
Figure 1 shows that as the variance of the clusters in the Gaussian mixture decreases, the theoretical bound increasingly aligns with the average final cost achieved during training. In particular, in Figure 2, corresponding to the smallest variance, the bound closely tracks the final training loss and, for some runs, is even smaller than the attained cost.
Figure 2. (Average of) initial ($C_{init}$) and final ($C_{final}$) cost of randomly initialized neural network(s) trained to classify Gaussian mixture data with $Q$ classes, plotted against the bound (4.1) computed for each data set.
We summarize the experiment parameters, as well as the hyperparameters used in the training of the neural networks via stochastic gradient descent, in Table 1. Code is available at https://github.com/patriciaewald/bounds.
Hyperparameter Notation Value
Input dimension $M$ $10$
Samples per class $N_{j}$ $100$
Number of classes $Q$ $\{10,\cdots,2\}$
Cluster standard deviation std $\{0.5,\cdots,0.1\}$
Random initializations $100$
Maximum number of epochs $1000$
Learning rate $0.05$
Momentum $0.9$
Plateau tolerance $10^{-4}$
Plateau patience (epochs) $30$
Table 1. Hyperparameters used in the Gaussian mixture and network training experiments.
5. Proof of Theorem 3.1
Let $R\in O(M)$ diagonalize $P,P^{\perp}$ (which are symmetric),
$P_{R}:=RPR^{T}\;\;\;,\;\;\;P^{\perp}_{R}:=RP^{\perp}R^{T}\,$, (5.1)
where $P_{R}$ has $Q$ diagonal entries of $1$, and all other components are 0, while $P^{\perp}_{R}$ has MQM-Q diagonal entries of $1$, and all other components are 0. This does not characterize $R$ uniquely, as we may compose $R$ with an arbitrary $R^{\prime}\in O(M)$ that leaves the ranges of $P,P^{\perp}$ invariant. Diagonality of $P_{R},P_{R}^{\perp}$ implies that the activation function, which acts component-wise as the ramp function (2.5), satisfies
$\sigma(P_{R}v+P^{\perp}_{R}v)=\sigma(P_{R}v)+\sigma(P^{\perp}_{R}v)$ (5.2)
for any $v\in{\mathbb{R}}^{M}$, and where of course, $P_{R}+P_{R}^{\perp}={\bf 1}_{M\times M}$.
We write
$b_{1}=P_{R}b_{1}+P_{R}^{\perp}b_{1}$ (5.3)
where we choose
$P_{R}b_{1}=\beta_{1}u_{Q}\,$, (5.4)
and we pick
$\beta_{1}\geq 2\rho\;\;\;,\;\;\rho:=\max_{j,i}|x_{0,j,i}|\,.$ (5.5)
This ensures that $RPx_{0,j,i}+P_{R}b_{1}\in B_{\rho}(\beta_{1}u_{Q})\subset P_{R}{\mathbb{R}}_{+}^{M}$ so that
$\sigma(RPx_{0,j,i}+P_{R}b_{1})=RPx_{0,j,i}+P_{R}b_{1}\;\;\;,\;\;\;j=1,\dots,Q\;\;,\;\forall\;i=1,\dots,N_{j}$ (5.6)
To construct the upper bound, we choose
$W_{1}=R\,.$ (5.7)
Then,
$X^{(1)}$ $=$ $\sigma(W_{1}X_{0}+B_{1})$ (5.8)
$=$ $\sigma((P_{R}+P_{R}^{\perp})RX_{0}+(P_{R}+P_{R}^{\perp})B_{1})$
$=$ $\sigma(P_{R}RX_{0}+P_{R}B_{1})+\sigma(P^{\perp}_{R}RX_{0}+P^{\perp}_{R}B_{1})$
$=$ $\sigma(RPX_{0}+P_{R}B_{1})+\sigma(RP^{\perp}\Delta X_{0}+P^{\perp}_{R}B_{1})\,.$
This is because by construction,
$P\overline{X_{0}}=\overline{X_{0}}\;\;\;,\;\;P^{\perp}\overline{X_{0}}=0\,.$ (5.9)
Moreover, we used
$P_{R}R=RP\;\;\;,\;\;P^{\perp}_{R}R=RP^{\perp}\,$, (5.10)
from (5.1).
Going to the next layer, we have
$X^{(2)}=W_{2}X^{(1)}+B_{2}\,.$ (5.11)
Our goal is to minimize the expression on the right hand side of
$\sqrt{N}{\mathcal{C}}[W_{j},b_{j}]$ $\leq$ $\|W_{2}\sigma(RPX_{0}+P_{R}B_{1})+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}}$ (5.12)
$\hskip 28.45274pt+\|W_{2}\sigma(RP^{\perp}\Delta X_{0}+P^{\perp}_{R}B_{1})\|_{{\mathcal{L}}^{2}}\,.$
We note that because ${\rm rank}(Y^{ext})=Q$, and
${\rm rank}(\sigma(\underbrace{RPX_{0}}_{=P_{R}RPX_{0}}+P_{R}B_{1}))\leq Q\,$, (5.13)
minimization of the first term on the r.h.s. of (5.12) requires that the latter has maximal rank $Q$. Due to our choice of $R$ and $P_{R}b_{1}$ which imply that (5.6) holds, we find that
$\sigma(RPX_{0}+P_{R}B_{1})$ $=$ $RPX_{0}+P_{R}B_{1}$ (5.14)
$=$ $P_{R}R(\overline{X_{0}}+P\Delta X_{0})+P_{R}B_{1}\,.$
To minimize the first term on the r.h.s. of (5.12), one might want to try to solve
$W_{2}S=Y^{ext}$ (5.15)
where
$S:=P_{R}R(\overline{X_{0}}+P\Delta X_{0})+P_{R}B_{1}\in{\mathbb{R}}^{M\times N}$ (5.16)
has rank $Q<M$ (due to $P_{R}$). However, this implies that the $QM$ unknown matrix components of $W_{2}\in{\mathbb{R}}^{Q\times M}$ are determined by $QN>QM$ equations. This is an overdetermined problem which generically has no solution. However, we can solve the problem to zeroth order in $\Delta X_{0}$,
$W_{2}S_{0}=Y^{ext}$ (5.17)
with
$S_{0}:=S|_{\Delta X_{0}=0}=P_{R}R\overline{X_{0}}=RP\overline{X_{0}}$ (5.18)
(recalling from (2.17) that $\overline{X_{0}}$ only contains $Q$ distinct column vectors $\overline{x_{0,j}}$ and their identical copies) and subsequently minimize the resulting expression in the first term on the r.h.s. of (5.12) with respect to $P_{R}B_{1}$.
This is equivalent to requiring that $W_{2}\in{\mathbb{R}}^{Q\times M}$ solves
$W_{2}RP\overline{X_{0}^{red}}=Y\,.$ (5.19)
To this end, we make the ansatz
$W_{2}=A(R\overline{X_{0}^{red}})^{T}$ (5.20)
for some $A\in{\mathbb{R}}^{Q\times Q}$. Then, recalling that $P\overline{X_{0}^{red}}=\overline{X_{0}^{red}}$,
$A(R\overline{X_{0}^{red}})^{T}(R\overline{X_{0}^{red}})=A(\overline{X_{0}^{red}})^{T}(\overline{X_{0}^{red}})=Y\,$, (5.21)
so that solving for $A$, we get
$W_{2}$ $=$ $Y((\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}})^{-1}(R\overline{X_{0}^{red}})^{T}$ (5.22)
$=$ $\widetilde{W}_{2}R^{T}\,$
with $\widetilde{W}_{2}$ as defined in (3.34). Hereby, $W_{2}$ is fully determined by $Y$ and $\overline{X_{0}^{red}}$ (as $\overline{X_{0}^{red}}$ determines $P$ and $R$).
Next, we can choose $P^{\perp}_{R}b_{1}$ in a suitable manner that
$\sigma(RP^{\perp}\Delta X_{0}+P^{\perp}_{R}B_{1})=0\,.$ (5.23)
This can be accomplished with
$P^{\perp}_{R}b_{1}=-\delta P^{\perp}_{R}u_{M}\in{\mathbb{R}}^{M}$ (5.24)
where $\delta$ is defined in (2.21), and
$u_{M}:=(1,1,\dots,1)^{T}\in{\mathbb{R}}^{M}\,.$ (5.25)
With
$E:=u_{M}u_{N}^{T}=[u_{M}\cdots u_{M}]\in{\mathbb{R}}^{M\times N}\,$, (5.26)
this ensures that the argument on the l.h.s. of (5.23), given by
$RP^{\perp}\Delta X_{0}+P^{\perp}_{R}B_{1}=P^{\perp}_{R}R\Delta X_{0}-\delta P^{\perp}_{R}E\,$, (5.27)
is a matrix all of whose components are $\leq 0$. Therefore, $\sigma$ maps it to 0.
Consequently, we find
$\sqrt{N}{\mathcal{C}}[W_{j},b_{j}]\leq\|W_{2}(RP\Delta X_{0}+P_{R}B_{1})+B_{2}\|_{{\mathcal{L}}^{2}}\,.$ (5.28)
To minimize the r.h.s., we may rewrite the bias
$b_{2}\rightarrow b_{2}-W_{2}P_{R}b_{1}$ (5.29)
so that it remains to minimize the r.h.s. of
$\sqrt{N}{\mathcal{C}}[W_{j},b_{j}]\leq\|W_{2}RP\Delta X_{0}+B_{2}\|_{{\mathcal{L}}^{2}}\,.$ (5.30)
with respect to
$B_{2}=[b_{2}\cdots b_{2}]=b_{2}u_{N}^{T}\;\;\in{\mathbb{R}}^{Q\times N}\,.$ (5.31)
We find
$0$ $=$ $\partial_{b_{2}}\|W_{2}RP\Delta X_{0}+\beta_{2}u_{N}^{T}\|_{{\mathcal{L}}^{2}}^{2}$ (5.32)
$=$ $\partial_{b_{2}}\Big(2{\rm Tr}(W_{2}RP\underbrace{\Delta X_{0}u_{N}}_{=0}b_{2}^{T})+{\rm Tr}(u_{N}\underbrace{b_{2}^{T}b_{2}}_{=|b_{2}|^{2}}u_{N}^{T})\Big)\,$,
using cyclicity of the trace. The first term inside the bracket on the second line is zero because for every $i\in\{1,\dots,M\}$,
$(\Delta X_{0}u_{N})_{i}$ $=$ $\sum_{\ell=1}^{N}[\Delta X_{0}]_{i\ell}$ (5.33)
$=$ $\sum_{j=1}^{Q}\sum_{\ell=1}^{N_{j}}(x_{0,j,\ell}-\overline{x_{0,j}})_{i}$
$=$ $0\,.$
The second term inside the bracket in (5.32) is proportional to $|b_{2}|^{2}$, so that its gradient is proportional to $b_{2}$, and vanishes for $b_{2}=0$. Hence, $B_{2}=0$, which means that
$B_{2}=-W_{2}P_{R}B_{1}$ (5.34)
for the original variable $b_{2}$ prior to the transformation (5.29).
We thus arrive at an upper bound
$\min_{W_{j},b_{j}}\sqrt{N}{\mathcal{C}}[W_{j},b_{j}]$ $\leq$ $\|\widetilde{W}_{2}P\Delta X_{0}\|_{{\mathcal{L}}^{2}}$ (5.35)
$=$ $\|Y{\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}\|_{{\mathcal{L}}^{2}}\,.$
This establishes the asserted upper bound in (3.15).
Next, we bound
$\|\widetilde{W}_{2}P\Delta X_{0}\|_{{\mathcal{L}}^{2}}$ $\leq$ $\|Y\|_{op}\|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}\|_{{\mathcal{L}}^{2}}\,.$ (5.36)
Let for notational brevity $A:={\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}$. We have
$\|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}\|_{{\mathcal{L}}^{2}}^{2}$ $=$ ${\rm Tr}(AA^{T})$ (5.37)
$=$ ${\rm Tr}(A^{T}A)$
$=$ $\sum_{\ell=1}^{Q}\sum_{i=1}^{N_{j}}|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,\ell,i}|^{2}$
$\leq$ $\Big(\sup_{i,\ell}\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,\ell,i}\right|^{2}\Big)\sum_{\ell=1}^{Q}\sum_{i=1}^{N_{\ell}}1$
$=$ $N\sup_{i,\ell}\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,\ell,i}\right|^{2}\,.$
Therefore,
$\min_{W_{j},b_{j}}\sqrt{N}{\mathcal{C}}[W_{j},b_{j}]$ $\leq$ $\|\widetilde{W}_{2}P\Delta X_{0}\|_{{\mathcal{L}}^{2}}$ (5.38)
$\leq$ $\sqrt{N}\;\|Y\|_{op}\sup_{i,\ell}\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,\ell,i}\right|$
$\leq$ $\sqrt{N}\;\|Y\|_{op}\delta_{P}\,$,
which proves (3.16). This concludes the proof of Theorem 3.1. ∎
6. Proof of Theorem 3.2
To prove the exact degenerate local minimum in the special case $M=Q$, we note that in this situation,
$P={\bf 1}_{Q\times Q}\;\;\;,\;P^{\perp}=0\,.$ (6.1)
In place of (5.12), the cost function is given by
${\mathcal{C}}_{{\mathcal{N}}}[W_{j},b_{j}]$ $=$ $\|W_{2}\sigma(W_{1}(\overline{X_{0}}+\Delta X_{0})+B_{1})+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}_{\mathcal{N}}}\,$, (6.2)
defined in (3.22), which is weighted by the inverse of the block diagonal matrix
${\mathcal{N}}={\rm diag}(N_{j}{\bf 1}_{N_{j}\times N_{j}}\,|\,j=1,\dots,Q)\;\in{\mathbb{R}}^{N\times N}\,.$ (6.3)
Using a similar choice for $B_{1}$ as in (5.4), we may assume that the matrix components of $W_{1}(\overline{X_{0}}+\Delta X_{0})+B_{1}$ are non-negative. To be precise, we let
$b_{1}=\beta_{1}u_{Q}$ (6.4)
for some $\beta_{1}\geq 0$ large enough that
$x_{0,j,i}+\beta_{1}u_{Q}\in{\mathbb{R}}_{+}^{Q}$ (6.5)
for all $j=1,\dots,Q$, $i=1,\dots,N_{j}$. This is achieved with
$\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|\,.$ (6.6)
Hence, choosing
$W_{1}={\bf 1}_{Q\times Q}\,$, (6.7)
we find
$\sigma(W_{1}X_{0}+B_{1})=\overline{X_{0}}+\Delta X_{0}+B_{1}\,.$ (6.8)
Thus, 6.2 implies
$\min_{W_{j},b_{j}}{\mathcal{C}}_{{\mathcal{N}}}[W_{j},b_{j}]\leq\|W_{2}(\overline{X_{0}}+\Delta X_{0})+W_{2}B_{1}+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}_{\mathcal{N}}}\,.$ (6.9)
To minimize the r.h.s., we note that since $W_{2}$ and $B_{2}$ are both unknown, we can redefine $W_{2}\rightarrow W_{2}$, and we can bias
$b_{2}\rightarrow b_{2}-W_{2}b_{1}\,.$ (6.10)
Therefore, we find
$\min_{W_{j},b_{j}}{\mathcal{C}}_{{\mathcal{N}}}[W_{j},b_{j}]\leq\min_{W_{2},b_{2}}\|W_{2}(\overline{X_{0}}+\Delta X_{0})+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}_{\mathcal{N}}}\,.$ (6.11)
Minimizing with respect to $W_{2}\in{\mathbb{R}}^{Q\times Q}$ is a least squares problem whose solution has the form
$(W_{2}(\overline{X_{0}}+\Delta X_{0})-Y^{ext}_{b_{2}}){\mathcal{P}}=0\,$, (6.12)
where
$Y^{ext}_{b_{2}}:=Y^{ext}-B_{2}\,$, (6.13)
${\mathcal{P}}^{T}$ is a projector onto the range of $X_{0}^{T}$, and remains to be determined. To solve (6.12), we require $W_{2}$ to satisfy
$W_{2}(\overline{X_{0}}+\Delta X_{0}){\mathcal{N}}^{-1}(\overline{X_{0}}+\Delta X_{0})^{T}$ (6.14)
$\hskip 56.9055pt=Y^{ext}_{b_{2}}{\mathcal{N}}^{-1}(\overline{X_{0}}+\Delta X_{0})^{T}\;\;\in{\mathbb{R}}^{Q\times Q}\,.$
Because
$\overline{X_{0}}$ $=$ $[\overline{x_{0,1}}u_{N_{1}}^{T}\cdots\overline{x_{0,Q}}u_{N_{Q}}^{T}]\,$,
$Y^{ext}_{b_{2}}$ $=$ $[(y_{1}-b_{2})u_{N_{1}}^{T}\cdots(y_{Q}-b_{2})u_{N_{Q}}^{T}]\,$, (6.15)
where
$u_{N_{j}}:=(1,1,\dots,1)^{T}\;\;\in{\mathbb{R}}^{N_{j}}\,$, (6.16)
we find that
$\Delta X_{0}{\mathcal{N}}^{-1}(\overline{X_{0}})^{T}$ $=$ $\sum_{j=1}^{Q}\frac{1}{N_{j}}\Delta X_{0,j}u_{N_{j}}\overline{x_{0,j}}^{T}$ (6.17)
$=$ $\sum_{j=1}^{Q}\underbrace{\Big(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\Delta x_{0,j,i}\Big)}_{=0}\overline{x_{0,j}}^{T}$
$=$ $0\,.$
For the same reason,
$\overline{X_{0}}{\mathcal{N}}^{-1}\Delta X_{0}^{T}=0\;\;\;,\;Y^{ext}{\mathcal{N}}^{-1}\Delta X_{0}^{T}=0\,.$ (6.18)
Moreover,
$Y^{ext}_{b_{2}}{\mathcal{N}}^{-1}\overline{X_{0}}^{T}$ $=$ $\sum_{j=1}^{Q}\frac{1}{N_{j}}(y_{j}-b_{2})\underbrace{u_{N_{j}}^{T}u_{N_{j}}}_{=N_{j}}\overline{x_{0,j}}^{T}$ (6.19)
$=$ $Y_{b_{2}}\overline{X_{0}^{red}}^{T}\,$,
where
$Y_{b_{2}}=[(y_{1}-b_{2})\cdots(y_{Q}-b_{2})]\;\;\in{\mathbb{R}}^{Q\times Q}$ (6.20)
and
$\overline{X_{0}}{\mathcal{N}}^{-1}\overline{X_{0}}^{T}=\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}\,.$ (6.21)
Therefore, (6.14) reduces to
$W_{2}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}+\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})=Y_{b_{2}}\overline{X_{0}^{red}}^{T}\,$, (6.22)
and invertibility of $\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}\in{\mathbb{R}}^{Q\times Q}$ implies invertibility of the matrix in brackets on the l.h.s. of (6.22), because
$D_{2}[X_{0}]:=\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T}=\sum_{j=1}^{Q}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\Delta x_{0,j,i}\Delta x_{0,j,i}^{T}\;\;\in{\mathbb{R}}^{Q\times Q}\,$ (6.23)
is a non-negative operator. Hence,
$W_{2}=Y_{b_{2}}\overline{X_{0}^{red}}^{T}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}+\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})^{-1}\,$, (6.24)
so that
$W_{2}X_{0}=Y^{ext}_{b_{2}}{\mathcal{P}}\,.$ (6.25)
Here,
${\mathcal{P}}:={\mathcal{N}}^{-1}X_{0}^{T}(X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}\in{\mathbb{R}}^{N\times N}$ (6.26)
is a rank $Q$ projector ${\mathcal{P}}^{2}={\mathcal{P}}$ which is orthogonal with respect to the inner product on ${\mathbb{R}}^{N}$ defined by $\mathcal{N}$}, i.e., ${\mathcal{P}}^{T}{\mathcal{N}}={\mathcal{N}}{\mathcal{P}}$. ${\mathcal{P}}^{T}$ is the projector onto the range of $X_{0}^{T}$. We note that with respect to the inner product defined with ${\mathcal{N}}^{-1}$ (which we are using here), we have ${\mathcal{N}}^{-1}{\mathcal{P}}^{T}={\mathcal{P}}{\mathcal{N}}^{-1}$.
Next, in order to control $W_{2}X_{0}-Y^{ext}_{b_{2}}=-Y^{ext}_{b_{2}}{\mathcal{P}}^{\perp}$, we observe that
$Y^{ext}_{b_{2}}{\mathcal{P}}$ $=$ $Y^{ext}_{b_{2}}{\mathcal{N}}^{-1}X_{0}^{T}(X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}$
$=$ $Y_{b_{2}}\overline{X_{0}^{red}}^{T}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}+D_{2}[X_{0}])^{-1}X_{0}$
$=$ $Y_{b_{2}}\overline{X_{0}^{red}}^{T}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T})^{-1}X_{0}$
$-Y_{b_{2}}\overline{X_{0}^{red}}^{T}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T})^{-1}D_{2}[X_{0}](X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}$
$=$ $Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}X_{0}$
$-Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}D_{2}[X_{0}](X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}$
where we have used the matrix identity
$(A+B)^{-1}$ $=$ $A^{-1}-A^{-1}B(A+B)^{-1}$ (6.28)
$=$ $A^{-1}-(A+B)^{-1}BA^{-1}$
for $A$, $A+B$ invertible. We observe that
$Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}X_{0}$ $=$ $Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}\overline{X_{0}}+Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}\Delta{X_{0}}$ (6.29)
$=$ $Y^{ext}_{b_{2}}+Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}\Delta{X_{0}}\,.$
To pass to the last line, we used that $(\overline{X_{0}^{red}})^{-1}\overline{X_{0}}=\,[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]$ where $\{e_{j}\in{\mathbb{R}}^{Q}\,|\,j=1,\dots,Q\}$ are the unit basis vectors.
We therefore conclude that, due to ${\mathcal{P}}^{\perp}={\bf 1}-{\mathcal{P}}$,
$Y^{ext}_{b_{2}}{\mathcal{P}}^{\perp}$ $=$ $-Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}\Delta{X_{0}}+{\mathcal{R}}$ (6.30)
where
${\mathcal{R}}$ $:=$ $Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}D_{2}[X_{0}](X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}$ (6.31)
satisfies
${\mathcal{R}}{\mathcal{P}}={\mathcal{R}}\,$, (6.32)
as can be easily verified.
At this point, we have found that
$\min_{W_{j},b_{j}}{\mathcal{C}}_{{\mathcal{N}}}[W_{j},b_{j}]\leq\min_{b_{2}}\|Y^{ext}_{b_{2}}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}\,.$ (6.33)
To minimize with respect to $b_{2}$, we set
$0$ $=$ $\partial_{b_{2}}\|Y^{ext}_{b_{2}}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}$ (6.34)
$=$ $\partial_{b_{2}}\Big(-2{\rm Tr}\Big(Y^{ext}{\mathcal{P}}^{\perp}{\mathcal{N}}^{-1}({\mathcal{P}}^{\perp})^{T}u_{N}b_{2}^{T}\Big)+{\rm Tr}\big(b_{2}u_{N}^{T}{\mathcal{P}}^{\perp}{\mathcal{N}}^{-1}({\mathcal{P}}^{\perp})^{T}u_{N}b_{2}^{T}\big)\Big)$
$=$ $\partial_{b_{2}}\Big(-2{\rm Tr}\Big(b_{2}^{T}Y^{ext}{\mathcal{P}}^{\perp}{\mathcal{N}}^{-1}({\mathcal{P}}^{\perp})^{T}u_{N}\Big)+|{\mathcal{N}}^{-1/2}({\mathcal{P}}^{\perp})^{T}u_{N}|^{2}|b_{2}|^{2}\Big)$
$=$ $\partial_{b_{2}}\Big(-2{\rm Tr}\Big(b_{2}^{T}Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}\underbrace{\Delta{X_{0}}{\mathcal{N}}^{-1}u_{N}}_{=0}\Big)\;+\;{\rm Tr}\Big(b_{2}^{T}\underbrace{{\mathcal{R}}{\mathcal{N}}^{-1}({\mathcal{P}}^{\perp})^{T}}_{={\mathcal{R}}{\mathcal{P}}{\mathcal{N}}^{-1}{\mathcal{P}}^{\perp}=0}u_{N}\Big)$
$\hskip 56.9055pt+|{\mathcal{N}}^{-1/2}({\mathcal{P}}^{\perp})^{T}u_{N}|^{2}|b_{2}|^{2}\Big)$
using ${\mathcal{R}}={\mathcal{R}}{\mathcal{P}}$ from (6.32) so that ${\mathcal{R}}{\mathcal{N}}^{-1}{\mathcal{P}}^{\perp}={\mathcal{R}}{\mathcal{P}}{\mathcal{N}}^{-1}{\mathcal{P}}^{\perp}={\mathcal{R}}{\mathcal{P}}{\mathcal{P}}^{\perp}{\mathcal{N}}^{-1}=0$, and recalling $u_{N}=(1,1,\dots,1)^{T}\in{\mathbb{R}}^{N}$. This implies that
$b_{2}=0\,.$ (6.35)
We note that this corresponds to
$b_{2}=-W_{2}b_{1}$ (6.36)
prior to the bias (6.10), and hence arrive at (3.29), as claimed.
Next, we derive a simplified expression for $\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}$. For notational convenience, let
$\Delta_{1}^{rel}$ $:=$ $(\overline{X_{0}^{red}})^{-1}\Delta X_{0}$
$\Delta_{2}^{rel}$ $:=$ $(\overline{X_{0}^{red}})^{-1}D_{2}[X_{0}](\overline{X_{0}^{red}})^{-T}$ (6.37)
$=$ $\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}$
where we recall (6.23). Then, we obtain (with $b_{2}=0$)
${\mathcal{R}}$ $=$ $Y\Delta_{2}^{rel}\big(1+\Delta_{2}^{rel}\big)^{-1}(\overline{X_{0}^{red}})^{-1}X_{0}$ (6.38)
$=$ $Y\Delta_{2}^{rel}\big(1+\Delta_{2}^{rel}\big)^{-1}\Big([e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]+\Delta_{1}^{rel}\Big)$
where we used $X_{0}{\mathcal{N}}^{-1}X_{0}^{T}=\overline{X_{0}^{red}}(1+\Delta_{2}^{rel})(\overline{X_{0}^{red}})^{T}$ in (6.31) to obtain the first line. We find that
$-Y^{ext}{\mathcal{P}}^{\perp}$ (6.39)
$=$ $Y\Delta_{1}^{rel}-Y\Delta_{2}^{rel}\big(1+\Delta_{2}^{rel}\big)^{-1}\Big([e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]+\Delta_{1}^{rel}\Big)$
$=$ $Y(1-\Delta_{2}^{rel}\big(1+\Delta_{2}^{rel}\big)^{-1})\Delta_{1}^{rel}-Y\Delta_{2}^{rel}\big(1+\Delta_{2}^{rel}\big)^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]$
$=$ $Y(1+\Delta_{2}^{rel})^{-1}\Delta_{1}^{rel}-Y\Delta_{2}^{rel}\big(1+\Delta_{2}^{rel}\big)^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]\,.$
Therefore, we obtain that
$\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}$ $=$ $(I)+(II)+(III)$ (6.40)
where
$(I)$ $:=$ ${\rm Tr}\Big(Y(1+\Delta_{2}^{rel})^{-1}\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}(1+\Delta_{2}^{rel})^{-T}Y^{T}\Big)$ (6.41)
$=$ ${\rm Tr}\Big(Y(1+\Delta_{2}^{rel})^{-1}\Delta_{2}^{rel}(1+\Delta_{2}^{rel})^{-1}Y^{T}\Big)$
where we note that $\Delta_{2}^{rel}=(\Delta_{2}^{rel})^{T}\in{\mathbb{R}}^{Q\times Q}$ is symmetric. Moreover,
$(II)$ $:=$ $-2{\rm Tr}\Big(Y(1+\Delta_{2}^{rel})^{-1}\Delta_{1}^{rel}{\mathcal{N}}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]^{T}\big(1+\Delta_{2}^{rel}\big)^{-1}\Delta_{2}^{rel}Y^{T}\Big)$ (6.42)
$=$ $-\sum_{j=1}^{Q}2{\rm Tr}\Big(Y(1+\Delta_{2}^{rel})^{-1}(\overline{X_{0}^{red}})^{-1}\underbrace{\frac{1}{N_{j}}\Delta X_{0,j}u_{N_{j}}e_{j}^{T}}_{=0}\big(1+\Delta_{2}^{rel}\big)^{-1}\Delta_{2}^{rel}Y^{T}\Big)$
$=$ $0$
where we used that $\frac{1}{N_{j}}\Delta X_{0,j}u_{N_{j}}=\overline{\Delta X_{0,j}}=0$ for all $j=1,\dots,Q$. Furthermore,
$(III)$ $:=$ ${\rm Tr}\Big(Y\Delta_{2}^{rel}\big(1+\Delta_{2}^{rel}\big)^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]{\mathcal{N}}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]^{T}$ (6.43)
$\hskip 56.9055pt\big(1+\Delta_{2}^{rel}\big)^{-1}\Delta_{2}^{rel}Y^{T}\Big)$
$=$ ${\rm Tr}\Big(Y\Delta_{2}^{rel}\big(1+\Delta_{2}^{rel}\big)^{-1}\sum_{j=1}^{Q}\underbrace{\frac{1}{N_{j}}u_{N_{j}}^{T}u_{N_{j}}}_{=1}e_{j}e_{j}^{T}\big(1+\Delta_{2}^{rel}\big)^{-1}\Delta_{2}^{rel}Y^{T}\Big)$
$=$ ${\rm Tr}\Big(Y\Delta_{2}^{rel}\big(1+\Delta_{2}^{rel}\big)^{-1}\big(1+\Delta_{2}^{rel}\big)^{-1}\Delta_{2}^{rel}Y^{T}\Big)$
$=$ ${\rm Tr}\Big(Y\big(1+\Delta_{2}^{rel}\big)^{-1}(\Delta_{2}^{rel})^{2}\big(1+\Delta_{2}^{rel}\big)^{-1}Y^{T}\Big)$
using $\sum_{j=1}^{Q}e_{j}e_{j}^{T}={\bf 1}_{Q\times Q}$, and commutativity of $\Delta_{2}^{rel}$ and $(1+\Delta_{2}^{rel})^{-1}$.
We conclude that
$\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}$ $=$ $(I)+(III)$ (6.44)
$=$ ${\rm Tr}\Big(Y\big(1+\Delta_{2}^{rel}\big)^{-1}\Delta_{2}^{rel}(1+\Delta_{2}^{rel})\big(1+\Delta_{2}^{rel}\big)^{-1}Y^{T}\Big)$
$=$ ${\rm Tr}\Big(Y\Delta_{2}^{rel}\big(1+\Delta_{2}^{rel}\big)^{-1}Y^{T}\Big)$
$=$ ${\rm Tr}\Big(Y|\Delta_{2}^{rel}|^{\frac{1}{2}}\big(1+\Delta_{2}^{rel}\big)^{-1}|\Delta_{2}^{rel}|^{\frac{1}{2}}Y^{T}\Big)$
$\leq$ $\|\big(1+\Delta_{2}^{rel}\big)^{-1}\|_{op}\;{\rm Tr}\Big(Y\Delta_{2}^{rel}Y^{T}\Big)$
since the matrix $\Delta_{2}^{rel}\geq 0$ is positive semidefinite. We note that
${\rm Tr}\Big(Y\Delta_{2}^{rel}Y^{T}\Big)$ $=$ ${\rm Tr}\Big(Y\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}Y^{T}\Big)$ (6.45)
$=$ $\|Y(\overline{X_{0}^{red}})^{-1}\Delta{X_{0}}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}\,.$
Moreover, letting $\lambda_{-}:=\inf{\rm spec}(\Delta_{2}^{rel})$ and $\lambda_{+}:=\sup{\rm spec}(\Delta_{2}^{rel})$ denote the smallest and largest eigenvalue of the non-negative, symmetric matrix $\Delta_{2}^{rel}\geq 0$,
$0\leq\lambda_{-}\leq\lambda_{+}=\|\Delta_{2}^{rel}\|_{op}=\|\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}\|_{op}\leq C\delta_{P}^{2}\,.$ (6.46)
Therefore,
$\|\big(1+\Delta_{2}^{rel}\big)^{-1}\|_{op}=(1+\lambda_{-})^{-1}\leq 1-C_{0}^{\prime}\delta_{P}^{2}$ (6.47)
for a constant $C_{0}^{\prime}=0$ if ${\rm rank}(\Delta_{2}^{rel})<Q$ whereby $\lambda_{-}=0$, and $C_{0}^{\prime}>0$ if ${\rm rank}(\Delta_{2}^{rel})=Q$.
We find the following bound from (6.44),
$\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}$ $\leq$ $\sqrt{\|\big(1+\Delta_{2}^{rel}\big)^{-1}\|_{op}\;{\rm Tr}\Big(Y\Delta_{2}^{rel}Y^{T}\Big)}$ (6.48)
$\leq$ $(1-C_{0}\delta_{P}^{2})\;\|Y(\overline{X_{0}^{red}})^{-1}\Delta{X_{0}}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}\,$,
as claimed, where $C_{0}$ is proportional to $C_{0}^{\prime}$.
Finally, we use that
$\widetilde{W}_{2}$ $=$ $Y\overline{X_{0}^{red}}^{T}(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T})^{-1}$ (6.49)
$=$ $Y(\overline{X_{0}^{red}})^{-1}\,$
agrees with (3.34) because $\overline{X_{0}^{red}}$ is invertible, and hence $(\overline{X_{0}^{red}})^{-1}={\rm Pen}[\overline{X_{0}^{red}}]$. Hence, we observe that (6.24) implies
$W_{2}$ $=$ $\widetilde{W}_{2}-\widetilde{W}_{2}(\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})(\overline{X_{0}^{red}}\,\overline{X_{0}^{red}}^{T}+\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})^{-1}$ (6.50)
$=$ $\widetilde{W}_{2}-Y(\overline{X_{0}^{red}})^{-1}(\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})(\overline{X_{0}^{red}})^{-T}$
$\hskip 56.9055pt\Big({\bf 1}+(\overline{X_{0}^{red}})^{-1}(\Delta X_{0}{\mathcal{N}}^{-1}\Delta X_{0}^{T})(\overline{X_{0}^{red}})^{-T}\Big)^{-1}(\overline{X_{0}^{red}})^{-1}$
$=$ $\widetilde{W}_{2}-Y\Delta_{2}^{rel}({\bf 1}+\Delta_{2}^{rel})^{-1}(\overline{X_{0}^{red}})^{-1}$
Therefore, in operator norm,
$\|W_{2}-\widetilde{W}_{2}\|_{op}$ $\leq$ $\|Y\|_{op}\|\Delta_{2}^{rel}\|_{op}\|({\bf 1}+\Delta_{2}^{rel})^{-1}\|_{op}\|(\overline{X_{0}^{red}})^{-1}\|_{op}$ (6.51)
$\leq$ $C\delta_{P}^{2}$
where we used $\|Y\|_{op}<C$, $\|\Delta_{2}^{rel}\|_{op}<C\delta_{P}^{2}$, $\|({\bf 1}+\Delta_{2}^{rel})^{-1}\|_{op}\leq 1$, $\|(\overline{X_{0}^{red}})^{-1}\|_{op}<C$ to obtain the last line.
We note that when $M>Q$, it follows that $\overline{X_{0}^{red}}(\overline{X_{0}^{red}})^{T}$ is an $M\times M$ matrix of rank $Q$, and is hence not invertible, so that the above arguments do not apply.
From here on, we denote the weights and biases determined above in (6.7), (6.24) and in (6.4), (6.36), by $W_{1}^{*},W_{2}^{*}$ and $b_{1}^{*},b_{2}^{*}$, respectively. To prove that ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is a local minimum of the cost function ${\mathcal{C}}_{{\mathcal{N}}}$, we observe that for
$\rho:=\max_{j,i}|x_{0,j,i}|$ (6.52)
and
$\beta_{1}\geq 2\rho\;\;\;,\;\;b_{1}^{*}=\beta_{1}u_{Q}\,$, (6.53)
we have
$W_{1}^{*}x_{0,j,i}+b_{1}^{*}$ $=$ $x_{0,j,i}+b_{1}^{*}$ (6.54)
$\in$ $B_{\rho}(\beta_{1}u_{Q})\subset{\mathbb{R}}_{+}^{Q}\,.$
$\forall j=1,\dots,Q$ and $\forall\;i=1,\dots,N_{j}$. That is, all training input vectors are contained in a ball of radius $\rho$ centered at a point on the diagonal $\beta_{1}u_{Q}\in{\mathbb{R}}^{Q}$ with $\beta_{1}\geq 2\rho$. This ensures that the coordinates of all $W_{1}^{*}x_{0,j,i}+b_{1}^{*}$ are strictly positive, and therefore,
$\sigma(W_{1}^{*}X_{0}+B_{1}^{*})=W_{1}^{*}X_{0}+B_{1}^{*}\,.$ (6.55)
For $\epsilon>0$ small, we consider an infinitesimal translation
$b_{1}^{*}\rightarrow b_{1}^{*}+\widetilde{b}_{1}$ (6.56)
combined with an infinitesimal transformation
$W_{1}^{*}\rightarrow W_{1}^{*}\widetilde{W}_{1}$ (6.57)
where $|\widetilde{b}_{1}|<\epsilon$ and $\|\widetilde{W}_{1}-{\bf 1}_{Q\times Q}\|_{op}<\epsilon$. Then, we have that
$W_{1}^{*}\widetilde{W}_{1}x_{0,j,i}+b_{1}^{*}+\widetilde{b}_{1}$ (6.58)
$=$ $\underbrace{W_{1}^{*}x_{0,j,i}+b_{1}^{*}}_{\in B_{\rho}(\beta_{1}u_{Q})}+W_{1}^{*}(\widetilde{W}_{1}-{\bf 1}_{Q\times Q})x_{0,j,i}+\widetilde{b}_{1}$
where
$|W_{1}^{*}(\widetilde{W}_{1}-{\bf 1}_{Q\times Q})x_{0,j,i}+\widetilde{b}_{1}|$ $\leq$ $\|W_{1}^{*}\|_{op}\|\widetilde{W}_{1}-{\bf 1}_{Q\times Q}\|_{op}|x_{0,j,i}|+|\widetilde{b}_{1}|$ (6.59)
$\leq$ $\|W_{1}^{*}\|_{op}\;\epsilon\;\max_{j,i}|x_{0,j,i}|+\epsilon$
$\leq$ $C\epsilon$
for some constant $C$ independent of $j,i$. Therefore, we conclude that (6.58) is contained in $B_{\rho+C\epsilon}(\beta_{1}u_{Q})$, which lies in the positive sector ${\mathbb{R}}_{+}^{Q}\subset{\mathbb{R}}^{Q}$ for sufficiently small $\epsilon>0$. But this implies that
$\sigma(W_{1}^{*}\widetilde{W}_{1}X_{0}+B_{1}^{*}+\widetilde{B}_{1})=W_{1}^{*}\widetilde{W}_{1}X_{0}+B_{1}^{*}+\widetilde{B}_{1}\,.$ (6.60)
Thus, all arguments leading to the result in (3.25) stating that
${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]=\|Y^{ext}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}\,$, (6.61)
for the choice of weights and biases $W_{i}^{*},b_{i}^{*}$, equally apply to the infinitesimally deformed $W_{1}^{*}\widetilde{W}_{1}$, $b_{1}^{*}+\widetilde{b}_{1}$ and the corresponding expressions for $W_{2}^{*}\widetilde{W}_{2}$, $b_{2}^{*}+\widetilde{b}_{2}$.
We now make the crucial observation that the right hand side of (6.61) does not depend on $W_{i}^{*},b_{i}^{*}$, as $\mathcal{P}$$^{\perp}$ only depends on the training inputs $X_{0}$; see (6.26).
Therefore,
${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]={\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*}\widetilde{W}_{i},b_{i}^{*}+\widetilde{b}_{i}]\,$, (6.62)
and the variation in the perturbations $\widetilde{W}_{i},\widetilde{b}_{i}$ is zero. This implies that ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is a local minimum of the cost function, as it was obtained from a least square minimization problem. In particular, it is degenerate, and by repeating the above argument, one concludes that it has the same value for all weights and biases $W_{i},b_{i}$ which allow for the condition (6.55) to be satisfied.
Finally, we note that the reparametrization of the training inputs $X_{0}\rightarrow KX_{0}$, for any arbitrary $K\in GL(Q)$, induces
$\overline{X_{0}}\rightarrow K\overline{X_{0}}\;\;\;,\;\;\Delta X_{0}\rightarrow K\Delta X_{0}\,.$ (6.63)
In particular, this implies that
$\Delta_{1}^{rel}$ $=$ $(\overline{X_{0}^{red}})^{-1}\Delta X_{0}$ (6.64)
$\rightarrow$ $(K\overline{X_{0}^{red}})^{-1}K\Delta X_{0}$
$=$ $(\overline{X_{0}^{red}})^{-1}K^{-1}K\Delta X_{0}$
$=$ $\Delta_{1}^{rel}$
is invariant under this reparametrization, and hence,
$\Delta_{2}^{rel}$ $=$ $\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}$ (6.65)
also is. From the third and fourth lines in (6.44) follows that ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is a function only of $\Delta_{2}^{rel}$. We note that the reparametrization invariance of $\mathcal{P}$} (and thus of $\mathcal{P}$$^{\perp}$) can also be directly verified from (6.26) by inspection. This implies that ${\mathcal{C}}_{{\mathcal{N}}}[W_{i}^{*},b_{i}^{*}]$ is invariant under reparametrizations of the training inputs $X_{0}\rightarrow KX_{0}$, for all $K\in GL(Q)$.
This concludes the proof of Theorem 3.2. ∎
7. Proof of Theorem 3.3
Here, we prove Theorem 3.3, and elucidate the geometric meaning of the constructively trained network obtained in Theorem 3.1. Given a test input $x\in{\mathbb{R}}^{M}$, the constructively trained network metrizes the linear subspace ${\rm range}(P)$ of the input space ${\mathbb{R}}^{M}$, and determines to which equivalence class of training inputs $Px$ is closest. To elucidate this interpretation, we observe that
$\widetilde{W}_{2}=Y((\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}})^{-1}(\overline{X_{0}^{red}})^{T}\in{\mathbb{R}}^{Q\times M}\,$ (7.1)
implies that
$\widetilde{W}_{2}P=\widetilde{W}_{2}\;\;\;{\rm and}\;W_{2}^{*}P_{R}=W_{2}^{*}\,$ (7.2)
and hence,
$\widetilde{W}_{2}\overline{X_{0}^{red}}=Y\,$, (7.3)
or in terms of column vectors,
$W_{2}^{*}RP\overline{x_{0,j}}=\widetilde{W}_{2}\overline{x_{0,j}}=y_{j}\,.$ (7.4)
Therefore,
${\mathcal{C}}_{j}[x]$ $=$ $|W_{2}^{*}(W_{1}^{*}x+b_{1}^{*})_{+}+b_{2}^{*}-y_{j}|$ (7.5)
$=$ $|W_{2}^{*}(Rx+b_{1}^{*})_{+}+b_{2}^{*}-y_{j}|$
$=$ $|W_{2}^{*}(RPx+RP^{\perp}x+P_{R}b_{1}^{*}+P^{\perp}_{R}b_{1}^{*})_{+}+b_{2}^{*}-y_{j}|$
$=$ $|W_{2}^{*}((P_{R}RPx+P_{R}b_{1}^{*})_{+}+(P^{\perp}_{R}RP^{\perp}x+P^{\perp}_{R}b_{1}^{*})_{+})+b_{2}^{*}-y_{j}|$
$=$ $|W_{2}^{*}(P_{R}RPx-y_{j})+W_{2}(P^{\perp}_{R}RP^{\perp}x-\delta P^{\perp}_{R}u)_{+}|$
$=$ $|W_{2}^{*}(P_{R}RP(x-\overline{x_{0,j}}))|$
where by construction of $P$, we have $P\overline{x_{0,j}}=\overline{x_{0,j}}$. Passing to the fourth line, we used $P_{R}RP=RP$, passing to the fifth line, we recalled $W_{2}^{*}P_{R}b_{1}^{*}+b_{2}^{*}=0$, and passing to the sixth line, we used (7.4). Moreover, we recall that $u$ is defined in (5.25). To pass to the sixth line, we also used that
$P^{\perp}_{R}(P^{\perp}_{R}RP^{\perp}x-\delta P^{\perp}_{R}u)_{+}=(P^{\perp}_{R}RP^{\perp}x-\delta P^{\perp}_{R}u)_{+}$ (7.6)
because the vector inside $(\cdot)_{+}$ only has components in the range of $P^{\perp}_{R}$, which is diagonal in the given coordinate system, and because $(\cdot)_{+}$ acts component-wise. Therefore, since also $W_{2}^{*}P_{R}=W_{2}^{*}$, we have
$W_{2}^{*}(P^{\perp}_{R}RP^{\perp}x-\delta P^{\perp}_{R}u)_{+}$ (7.7)
$=$ $W_{2}^{*}P_{R}P^{\perp}_{R}(P^{\perp}_{R}RP^{\perp}x-\delta P^{\perp}_{R}u)_{+}$
$=$ $0\,.$
With $W_{2}^{*}P_{R}RP=\widetilde{W}_{2}P$, (7.5) implies that
${\mathcal{C}}_{j}[x]=d_{\widetilde{W}_{2}}(Px,\overline{x_{0,j}})$ (7.8)
where
$d_{\widetilde{W}_{2}}(x,y):=|\widetilde{W}_{2}P(x-y)|\,$, (7.9)
for $x,y\in{\rm range}(P)$, defines a metric in the range of $P$, which is a $Q$-dimensional linear subspace of the input space ${\mathbb{R}}^{M}$. This is because the matrix $\widetilde{W}_{2}=\widetilde{W}_{2}P$ has full rank $Q$.
We can therefore reformulate the identification of an input $x\in{\mathbb{R}}^{M}$ with an output $y_{j^{*}}$ via the constructively trained shallow network as the solution of the metric minimization problem
$j^{*}={\rm argmin}_{j\in\{1,\dots,Q\}}(d_{\widetilde{W}_{2}}(Px,\overline{x_{0,j}}))$ (7.10)
in the range of $P$. ∎
Acknowledgments: T.C. gratefully acknowledges support by the NSF through the grant DMS-2009800, and the RTG Grant DMS-1840314 - Analysis of PDE. P.M.E. was supported by NSF grant DMS-2009800 through T.C.